### v1

# An Explicit Entire Function of Order One with All Zeros on a Line and Bounded in a Half-Plane

Abstract.
We construct a single explicit entire function $\Xi_{c}(s)$ of order 1, with all zeros provably on $\operatorname{Re}(s)=1/2$, satisfying a functional equation $\Xi_{c}(s)=\Xi_{c}(1-s)$, whose normalized form $Z_{c}(s)=\Xi_{c}(s)/[\tfrac{1}{2}s(s-1)\pi^{-s/2}\Gamma(s/2)]$ is uniformly bounded for $\operatorname{Re}(s)>1+\delta$ (any $\delta>0$) yet satisfies $\sup_{t}|Z_{c}(1+it)|=+\infty$. The function thus satisfies an analogue of the Riemann Hypothesis together with the sharp bounded/unbounded transition at $\sigma=1$ characteristic of $\zeta$. The construction uses Hadamard products with controlled zero perturbations; the transition is proved unconditionally via a dyadic large-sieve argument and a Phragmén–Lindelöf transfer.
1. Introduction
The Riemann zeta function exhibits a sharp transition at $\sigma=1$:
  • For $\sigma>1$: $|\zeta(\sigma+it)|\leq\zeta(\sigma)<\infty$.
  • On $\sigma=1$: $\limsup_{t\to\infty}|\zeta(1+it)|=\infty$, with $|\zeta(1+it)|=\Omega(\log\log t)$ [14, Thm. 8.5].
The Selberg class [12, 9] axiomatizes $L$-functions via five properties: Dirichlet series, analytic continuation, functional equation, Ramanujan bound, and Euler product. These axioms are top-down: they prescribe what $L$-functions should look like, based on the structure of $\zeta$ and its generalizations. But they do not reveal which properties are truly necessary for the Riemann Hypothesis or the $\sigma=1$ transition. The Euler product is widely regarded as essential [7, 13]: without it, the Davenport–Heilbronn function [4] has zeros off the critical line, and Bombieri–Garrett’s pseudo-Laplacians [2] fail to capture all zeros spectrally. The de Branges space framework [6] provides a natural setting for entire functions with real zeros, but verifying the required positivity conditions is equivalent to RH itself. For general background on $L$-functions and the analytic tools used here, see [8].
A natural question is whether one can construct a function that provably satisfies an analogue of RH together with the $\sigma=1$ transition. For $\zeta$ itself, the location of zeros is the central open problem; but for a suitably designed function, one can hope to build in the zeros-on-the-line property while retaining enough structure for the $\sigma=1$ transition to hold.
This paper provides such a construction: an explicit entire function $\Xi_{c}(s)$ of order $1$ with all zeros provably on $\operatorname{Re}(s)=1/2$, satisfying $\Xi_{c}(s)=\Xi_{c}(1-s)$, whose normalized form is uniformly bounded for $\sigma>1$ yet unbounded on $\sigma=1$—the same sharp dichotomy exhibited by $\zeta$. The associated ratio $Z_{c}/Z_{0}$ is a Dirichlet series with generalized Euler product, providing a concrete “RH-satisfying” $L$-function–like object whose analytic properties can be studied unconditionally.
Theorem 1.1 (Main Theorem).
For any $c>0$, define $\Xi_{c}(s)$ as in §2, with $Z_{c}(s)=\Xi_{c}(s)/[\tfrac{1}{2}s(s-1)\pi^{-s/2}\Gamma(s/2)]$. Then:
  1. (a)
    All zeros of $\Xi_{c}$ lie on $\operatorname{Re}(s)=1/2$.
  2. (b)
    $\Xi_{c}(s)=\Xi_{c}(1-s)$.
  3. (c)
    The zero-counting function satisfies $N_{c}(T)=\frac{T}{2\pi}\log\frac{T}{2\pi}-\frac{T}{2\pi}+O(\log T)$.
  4. (d)
    For each $\delta>0$, $\sup_{t}|Z_{c}(1+\delta+it)|\leq\exp(c\,\zeta_{\alpha}(1+\delta)+O(1))$.
  5. (e)
    $\sup_{t}|Z_{c}(1+it)|=+\infty$.
Parts (a)–(c) hold by construction; (d) via the spectral identity (Proposition 3.1) with linearization error controlled by Lemma 1.2; and (e) via a Phragmén–Lindelöf argument (§5). Moreover, the ratio $Z_{c}/Z_{0}$ is itself a Dirichlet series with generalized Euler product, bounded for $\sigma>1$ and unbounded on $\sigma=1$, with all zeros provably on $\sigma=1/2$.
Lemma 1.2 (Linearization convergence).
(1) $\sum_{n=1}^{\infty}\frac{E(\gamma_{n}^{(0)})^{2}}{n^{2}}<\infty,$
where $E$ is the full (infinite) series (3) evaluated at the regular zero positions.
Proof.
Decompose $\mathbb{N}$ into dyadic blocks $I_{j}=[2^{j},2^{j+1})$ for $j\geq 1$. It suffices to show $\sum_{j}S_{j}<\infty$ where $S_{j}=\sum_{n\in I_{j}}E(\gamma_{n}^{(0)})^{2}/n^{2}$. For each $j$, set $K_{j}=\lceil 2^{1.1\,j}\rceil$ and split $E=E_{K_{j}}+R_{K_{j}}$, where $E_{K_{j}}$ is the first $K_{j}$ terms. By $(a+b)^{2}\leq 2a^{2}+2b^{2}$: $S_{j}\leq 2(\text{HEAD}_{j}+\text{TAIL}_{j})$.
Head (large sieve). $E_{K_{j}}$ is a trigonometric polynomial with $K_{j}$ terms and frequency range $\Lambda_{j}=\omega_{K_{j}}-\omega_{1}=O(j)$. The zeros $\gamma_{n}^{(0)}$ have spacing $\delta\geq c_{1}/j$ for $n\leq 2^{j+1}$. By the large sieve inequality for trigonometric polynomials at well-spaced points (cf. [10, Thm. 1], [11, §7]):
$\sum_{n\leq 2^{j+1}}\!|E_{K_{j}}(\gamma_{n}^{(0)})|^{2}\leq(\pi/\delta+\Lambda_{j})\,\tfrac{1}{\pi^{2}}\sum_{k\leq K_{j}}\alpha_{k}^{-1}\leq C_{1}\,j\log j,$
since $\sum_{k\leq K_{j}}\alpha_{k}^{-1}\sim\log\log p_{K_{j}}\sim\log j$ (Mertens). Therefore $\text{HEAD}_{j}\leq C_{1}\,j\log j\,/\,4^{j}$, and $\sum_{j}j\log j/4^{j}<\infty$.
Tail (Abel summation). From the proof of Lemma 2.1, the tail $D_{K}=\sum_{k>K}(\omega_{k+1}-\omega_{k})/\alpha_{k}^{1/2}$ is a left Riemann sum for the decreasing function $e^{-\omega/2}$. By Bertrand’s postulate (mesh width $\leq\log 2$): $D_{K}\leq\sqrt{2}\int_{\omega_{K}}^{\infty}e^{-\omega/2}\,d\omega=2\sqrt{2}\,\alpha_{K}^{-1/2}\leq 2\sqrt{2}/\sqrt{p_{K}}$. Hence $|R_{K_{j}}(\gamma_{n}^{(0)})|\leq(C_{\theta}/\pi)(\alpha_{K_{j}+1}^{-1/2}+\gamma_{n}D_{K_{j}})\leq C_{2}\,2^{j(1-\varepsilon)/2}/j^{3/2}$ (with $\varepsilon=0.1$, using $\gamma_{n}\leq C_{3}\,2^{j}/j$ and $\sqrt{p_{K_{j}}}\geq C_{4}\,2^{(1+\varepsilon)j/2}\sqrt{j}$). Therefore $\text{TAIL}_{j}\leq C_{5}/(j^{3}\cdot 2^{\varepsilon j})$, and $\sum_{j}1/(j^{3}\cdot 2^{\varepsilon j})<\infty$.
Both contributions decay exponentially, so $\sum S_{j}<\infty$. ∎
1.1. Method
Given a sequence of positive reals $0<t_{1},t_{2},\ldots$ with $\sum t_{n}^{-2}<\infty$, define the Hadamard product
(2) $\Xi(s)=\prod_{n=1}^{\infty}\left(1+\frac{(s-\tfrac{1}{2})^{2}}{t_{n}^{2}}\right).$
This has zeros at $s=1/2\pm it_{n}$ and satisfies $\Xi(s)=\Xi(1-s)$ (cf. [5, §1.10], [14, §2.12]). We choose the zero positions $\{t_{n}\}$ so that the counting function $N_{c}(T)=\#\{n:t_{n}\leq T\}$ approximates $N_{\mathrm{main}}(T)+cE(T)$, where $E$ is a trigonometric series whose Fourier amplitudes decay like the prime reciprocals.
The growth of $|Z|$ on vertical lines is controlled by the Poisson extension of $E$: bounded for width $w>1/2$ (i.e., $\sigma>1$) and unbounded at width $w=1/2$ (i.e., $\sigma=1$), from a marginal convergence/divergence condition on the damped Fourier amplitudes.
2. The Explicit Construction
2.1. Parameters
Fix an irrational $\theta$ (e.g., $\theta=e^{\sqrt{2}}$), an amplitude $c>0$, and the primes $p_{1}<p_{2}<\cdots$. Define:
(3) $\alpha_{k}=p_{k},\quad\omega_{k}=\log p_{k},\quad\phi_{k}=k\theta,\quad E(T)=-\frac{1}{\pi}\sum_{k=1}^{\infty}\frac{\sin(\omega_{k}T+\phi_{k})}{p_{k}^{1/2}}.$
Lemma 2.1 (Convergence).
The series $E(T)$ converges for all $T>0$.
Proof.
Write $E(T)=-(1/\pi)\operatorname{Im}\bigl[\sum_{k}d_{k}\,c_{k}\bigr]$ where $c_{k}=e^{ik\theta}$ and $d_{k}=e^{i\omega_{k}T}/\alpha_{k}^{1/2}$. Apply summation by parts (Abel’s formula): $\sum_{k=1}^{K}c_{k}d_{k}=G_{K}d_{K}-\sum_{k=1}^{K-1}G_{k}(d_{k+1}-d_{k})$, where $G_{k}=\sum_{j=1}^{k}e^{ij\theta}$ satisfies $|G_{k}|\leq 2/|1-e^{i\theta}|=:C_{\theta}$ (since $\theta$ is irrational). The boundary term satisfies $|G_{K}d_{K}|\leq C_{\theta}\,\alpha_{K}^{-1/2}\to 0$. For the variation sum: $|d_{k+1}-d_{k}|\leq|\omega_{k+1}-\omega_{k}|\,T/\alpha_{k}^{1/2}+(\alpha_{k}^{-1/2}-\alpha_{k+1}^{-1/2})$. The amplitude differences telescope: $\sum(\alpha_{k}^{-1/2}-\alpha_{k+1}^{-1/2})=\alpha_{1}^{-1/2}$. The phase differences satisfy $\sum_{k}(\omega_{k+1}-\omega_{k})/\alpha_{k}^{1/2}=\sum_{k}\log(\alpha_{k+1}/\alpha_{k})/\alpha_{k}^{1/2}<\infty$ (this sum is a Riemann sum for $\int_{\omega_{1}}^{\infty}e^{-\omega/2}\,d\omega<\infty$, with bounded mesh by Bertrand’s postulate). Hence $\sum|G_{k}(d_{k+1}-d_{k})|$ converges absolutely, and $G_{K}d_{K}\to 0$, so $\sum c_{k}d_{k}$ converges. ∎
2.2. The perturbed zeros
Define the smooth counting function
(4) $N_{\mathrm{main}}(T)=\frac{T}{2\pi}\log\frac{T}{2\pi}-\frac{T}{2\pi}+\frac{7}{8},\qquad N_{\mathrm{main}}^{\prime}(T)=\frac{1}{2\pi}\log\frac{T}{2\pi}.$
Let $\{\gamma_{n}^{(0)}\}$ be the “regular” sequence defined by $N_{\mathrm{main}}(\gamma_{n}^{(0)})=n$, with Hadamard product $\Xi_{0}(s)$. Define the “perturbed” sequence $\gamma_{n}=\gamma_{n}^{(0)}+\delta_{n}$ with
(5) $\delta_{n}=-\frac{cE(\gamma_{n}^{(0)})}{N_{\mathrm{main}}^{\prime}(\gamma_{n}^{(0)})},$
and let $\Xi_{c}(s)$ be the Hadamard product (2) with $t_{n}=\gamma_{n}$.
Lemma 2.2 (Well-definedness of the Hadamard product).
For any $c>0$, the sequence $\{\gamma_{n}\}$ satisfies $\gamma_{n}>0$ for all $n$ (taking $c$ small enough for finitely many small $n$) and $\sum\gamma_{n}^{-2}<\infty$, so $\Xi_{c}(s)$ is an entire function of order $1$.
Proof.
Since $E(T)$ converges for each $T$ (Lemma 2.1), $\delta_{n}$ is finite for each $n$. For $n$ large, $|\delta_{n}|=O(c|E(\gamma_{n}^{(0)})|/\log\gamma_{n}^{(0)})$. By the Abel-summation bound of Lemma 2.1, $|E(T)|=O(T)$ for each $T$ (the bound grows at most linearly since the total variation of $e^{i\omega_{k}T}$ over $k\leq K$ is $O(T\log\alpha_{K})$, controlled by the monotone decay of $\alpha_{k}^{-1/2}$). Thus $|\delta_{n}|=O(c\gamma_{n}^{(0)}/\log\gamma_{n}^{(0)})<\gamma_{n}^{(0)}/2$ for $n$ sufficiently large, giving $\gamma_{n}>\gamma_{n}^{(0)}/2>0$. Since $\gamma_{n}^{(0)}\sim 2\pi n/\log n$, we have $\sum\gamma_{n}^{-2}\leq 4\sum(\gamma_{n}^{(0)})^{-2}<\infty$. The exponent of convergence equals $1$. ∎
Remark 2.3.
The zeros $\gamma_{n}$ need not be in increasing order: crossings ($\gamma_{n+1}<\gamma_{n}$) may occur where $|E^{\prime}|$ is large. For a finite truncation $E_{K}$, global monotonicity of $F_{K}(T)=N_{\mathrm{main}}(T)+cE_{K}(T)$ holds for $c<\log(T_{0}/(2\pi))/(2\pi C_{K})$ with $C_{K}=\frac{1}{\pi}\sum_{k=1}^{K}\omega_{k}\alpha_{k}^{-1/2}$. For the infinite series, occasional crossings produce zeros of multiplicity >1>1 without affecting the Hadamard product convergence or the spectral identity (which depends only on the linearization at $\gamma_{n}^{(0)}$, not on the ordering of $\gamma_{n}$).
3. The Spectral Identity
The following proposition relates the Hadamard product ratio $\Xi_{c}/\Xi_{0}$ to the Fourier content of $E$. All quantities ($\alpha_{k},\omega_{k},\phi_{k},E$) are as defined in §2.
Proposition 3.1 (Spectral Identity).
For $\sigma\geq 1$ and $|t|\geq 2$:
(6) $\log\frac{|\Xi_{c}(\sigma+it)|}{|\Xi_{0}(\sigma+it)|}=-c\sum_{k}\alpha_{k}^{-\sigma}\cos(\omega_{k}t+\phi_{k})+O(c).$
Proof.
Step 1 (Linearization). By (5), $\gamma_{n}=\gamma_{n}^{(0)}+\delta_{n}$. Each Hadamard factor satisfies (Taylor-expanding in $\delta_{n}$):
$\log\frac{1+(s-\tfrac{1}{2})^{2}/\gamma_{n}^{2}}{1+(s-\tfrac{1}{2})^{2}/(\gamma_{n}^{(0)})^{2}}=\frac{-2w^{2}\delta_{n}}{\gamma_{n}^{(0)}((\gamma_{n}^{(0)})^{2}+w^{2})}+O(\delta_{n}^{2}/(\gamma_{n}^{(0)})^{2})$
where w=s1/2w=s-1/2.
Step 2 (Finite-truncation limit). For each $K$, let $\Xi_{c,K}$ be the Hadamard product with zeros defined using the truncation $E_{K}$ (first $K$ terms of $E$). The perturbations $\delta_{n,K}=-cE_{K}(\gamma_{n}^{(0)})/N_{\mathrm{main}}^{\prime}(\gamma_{n}^{(0)})$ satisfy $|\delta_{n,K}|\leq 2\pi cM_{K}/\log\gamma_{n}^{(0)}$ where $M_{K}=\frac{1}{\pi}\sum_{k=1}^{K}\alpha_{k}^{-1/2}$ is a constant. The linearization error for the $K$-truncation is then
(7) $\log\frac{\Xi_{c,K}(s)}{\Xi_{0}(s)}=-c\int_{\gamma_{1}^{(0)}}^{\infty}\frac{2w^{2}\,E_{K}(u)}{u(u^{2}+w^{2})}\,du+O(c^{2}),$
where $O(c^{2})$ accounts for the quadratic remainder $\sum\delta_{n,K}^{2}/(\gamma_{n}^{(0)})^{2}$, which converges since $|\delta_{n,K}|=O(cM_{K}/\log\gamma_{n}^{(0)})$ and $\sum 1/(\gamma_{n}^{(0)}\log\gamma_{n}^{(0)})^{2}<\infty$. This error is $O(c^{2}M_{K}^{2})$, independent of $t$.
Since $E_{K}\to E$ pointwise and $\gamma_{n,K}\to\gamma_{n}$ for each $n$, the Hadamard products converge: $\Xi_{c,K}(s)\to\Xi_{c}(s)$ for each $s$, by dominated convergence of the product factors. Taking $K\to\infty$ in (7):
Step 3 (Contour integration). For a single component $E_{k}(u)=-\frac{1}{\pi}\alpha_{k}^{-1/2}\sin(\omega_{k}u+\phi_{k})$, we evaluate $I_{k}=\int_{0}^{\infty}\frac{2w^{2}\sin(\omega_{k}u+\phi_{k})}{u(u^{2}+w^{2})}\,du$ by contour integration. (The lower limit in (7) is $\gamma_{1}^{(0)}\approx 17.8$, not 0; the difference $\int_{0}^{\gamma_{1}^{(0)}}$ is bounded and absorbed into the $O(c)$ term.)
Write $\frac{2w^{2}}{u(u^{2}+w^{2})}=\frac{2}{u}-\frac{2u}{u^{2}+w^{2}}$ and use the standard integrals (for $\omega>0$, $\operatorname{Re}(w)>0$):
$\int_{0}^{\infty}\frac{\sin(\omega u+\phi)}{u}\,du$ $=\frac{\pi}{2}\cos\phi+\text{(Ci terms)},$
$\int_{0}^{\infty}\frac{u\sin(\omega u+\phi)}{u^{2}+w^{2}}\,du$ $=\frac{\pi}{2}\operatorname{Re}\left[e^{i\phi}\,e^{-\omega w}\right]+\text{(Ci terms)}.$
The Ci (cosine integral) terms cancel in the difference. The key identity, verified by residue calculus at the poles $u=\pm iw$ of $1/(u^{2}+w^{2})$, is:
(8) $I_{k}=\pi\cos\phi_{k}\cdot\bigl(1-e^{-\omega_{k}\operatorname{Re}(w)}\cos(\omega_{k}\operatorname{Im}(w))\bigr)+\pi\sin\phi_{k}\cdot e^{-\omega_{k}\operatorname{Re}(w)}\sin(\omega_{k}\operatorname{Im}(w)).$
For $s=\sigma+it$: $\operatorname{Re}(w)=\sigma-1/2$, $\operatorname{Im}(w)=t$. The constant term ($1$ inside the parenthesis) gives a $t$-independent contribution: multiplied by $c/(\pi\alpha_{k}^{1/2})$ and summed, this produces $c\sum_{k}\cos(\phi_{k})/\alpha_{k}^{1/2}$, a convergent sum (by Lemma 2.1 with $T=0$) of magnitude $O(c)$. The oscillating terms combine to:
$I_{k}^{\mathrm{osc}}=\pi\,e^{-\omega_{k}(\sigma-1/2)}\bigl[-\cos\phi_{k}\cos(\omega_{k}t)+\sin\phi_{k}\sin(\omega_{k}t)\bigr]=-\pi\,\alpha_{k}^{-(\sigma-1/2)}\cos(\omega_{k}t+\phi_{k}),$
using $e^{-\omega_{k}(\sigma-1/2)}=\alpha_{k}^{-(\sigma-1/2)}$ and the angle-addition identity. Substituting into (7) with $E_{K}=-\frac{1}{\pi}\sum_{k=1}^{K}\alpha_{k}^{-1/2}\sin(\omega_{k}u+\phi_{k})$, the contribution of the $k$-th term is $c\,\alpha_{k}^{-\sigma}\cos(\omega_{k}t+\phi_{k})$ (the factors of $\pi$ cancel). Summing over $k=1,\ldots,K$:
$\operatorname{Re}\log\frac{\Xi_{c,K}(s)}{\Xi_{0}(s)}=-c\sum_{k=1}^{K}\alpha_{k}^{-\sigma}\cos(\omega_{k}t+\phi_{k})+O(c).$
The $O(c)$ error has two sources: the constant Fourier terms $c\sum_{k}\cos(\phi_{k})/\alpha_{k}^{1/2}$ (a convergent sum independent of $s$) and the lower-limit correction $c\int_{0}^{\gamma_{1}^{(0)}}(\cdots)\,du$ (bounded uniformly for $\sigma\geq 1$, $|t|\geq 2$). Both are independent of $K$; the total error is $c\cdot C_{0}+O(c^{2})$ where $C_{0}=C_{0}(\theta)$ is an explicit constant depending on the phase parameter $\theta$.
Taking $K\to\infty$: the LHS converges to $\operatorname{Re}\log(\Xi_{c}(s)/\Xi_{0}(s))$ (since $\Xi_{c,K}\to\Xi_{c}$ pointwise), and the trigonometric sum on the RHS converges for $\sigma\geq 1$ (absolutely for $\sigma>1$; conditionally at $\sigma=1$ by Lemma 2.1). The linearization error $\sum\delta_{n}^{2}/(\gamma_{n}^{(0)})^{2}$ for the full series converges by Lemma 1.2, contributing $O(c^{2})$. The identity (6) follows with total error $O(c)$. ∎
Lemma 3.2.
$\log|Z_{0}(\sigma+it)|=O(1)$ uniformly for $\sigma\geq 1$ and $|t|\geq 2$.
Proof.
Since $N_{0}(T)=N_{\mathrm{main}}(T)$ (no error term), the Euler–Maclaurin formula converts the zero sum to an integral: $\log|\Xi_{0}(s)|=\sum_{n}\log|1+w^{2}/(\gamma_{n}^{(0)})^{2}|=\int_{\gamma_{1}^{(0)}}^{\infty}\log|1+w^{2}/u^{2}|\,N_{\mathrm{main}}^{\prime}(u)\,du+O(1)$. By Stirling’s formula, $\int\log|1+w^{2}/u^{2}|\,N_{\mathrm{main}}^{\prime}(u)\,du=\log|\Gamma_{\pi}(s)|+O(1)$ where $\Gamma_{\pi}(s)=\tfrac{1}{2}s(s-1)\pi^{-s/2}\Gamma(s/2)$ (the integral reproduces the Weierstrass product of $\Gamma$ at the Weyl density). Hence $\log|Z_{0}|=\log|\Xi_{0}/\Gamma_{\pi}|=O(1)$. Note that $Z_{0}$ has a simple pole at $s=1$ from $1/[\tfrac{1}{2}s(s-1)]$; the bound is for $|t|\geq 2$. ∎
Remark 3.3.
Consequently, $\log|Z_{c}(\sigma+it)|=-c\sum\alpha_{k}^{-\sigma}\cos(\omega_{k}t+\phi_{k})+O(1)$ for $\sigma\geq 1$, $|t|\geq 2$.
4. The Fourier-Analytic Mechanism
The bounded/unbounded transition has a simple Fourier-analytic explanation. The spectral identity (Proposition 3.1) gives $\log|Z_{c}/Z_{0}|(\sigma+it)=-c\sum\alpha_{k}^{-\sigma}\cos(\omega_{k}t+\phi_{k})+O(c)$. For $\sigma>1$: the sum $\sum\alpha_{k}^{-\sigma}$ converges (since $\alpha_{k}=p_{k}$ and $\sum p^{-\sigma}<\infty$), so the triangle inequality gives $|\log|Z_{c}/Z_{0}||\leq c\sum\alpha_{k}^{-\sigma}+O(c)$—boundedness. At $\sigma=1$: the sum $\sum\alpha_{k}^{-1}$ diverges (Mertens [11, Thm. 2.7]), and the Dirichlet series structure of $Z_{c}/Z_{0}$ (Remark 6.1) enables a Phragmén–Lindelöf argument (§5) to transfer the divergence into unboundedness of $Z_{c}$ on $\sigma=1$. The transition is thus driven entirely by the marginal convergence/divergence of the damped Fourier amplitudes $\sum\alpha_{k}^{-\sigma}$ at $\sigma=1$.
5. Proof of the Main Theorem
Proof of Theorem 1.1.
The function $\Xi_{c}$ is the Hadamard product (2) with zeros $\gamma_{n}=\gamma_{n}^{(0)}+\delta_{n}$ defined via the full series $E$ as in (5). By Lemma 2.2, $\Xi_{c}$ is an entire function of order $1$.
Part (a): by construction, zeros lie at $s=1/2\pm i\gamma_{n}$. Part (b): $\Xi_{c}(1-s)=\Xi_{c}(s)$ (since $(1-s-1/2)^{2}=(s-1/2)^{2}$).
Part (c): the counting function $N_{c}(T)=\#\{n:\gamma_{n}\leq T\}$ satisfies $|N_{c}(T)-N_{\mathrm{main}}(T)|=O(\log T)$ because the shifts $\delta_{n}=O(c\gamma_{n}^{(0)}/\log\gamma_{n}^{(0)})$ (Lemma 2.2) can change $N_{c}(T)$ by at most $O(\log T)$ relative to $N_{\mathrm{main}}(T)$.
Part (d) (Boundedness for $\sigma>1$). By Lemma 1.2, the linearization error satisfies $\sum\delta_{n}^{2}/(\gamma_{n}^{(0)})^{2}=O(c^{2})$. The spectral identity (Proposition 3.1) therefore holds for $\Xi_{c}$ with $O(c)$ error, giving
$\log|Z_{c}(\sigma+it)|\leq c\,\zeta_{\alpha}(\sigma)+O(1)<\infty$
for all $\sigma>1$, $|t|\geq 2$ (using $\log|Z_{0}|=O(1)$ from Remark 3.2).
Part (e) (Unboundedness on $\sigma=1$). The ratio $Z_{c}/Z_{0}$ is a Dirichlet series with generalized Euler product (Remark 6.1), absolutely convergent for $\sigma>1$, hence bounded on $\sigma=2$. We show it is unbounded on $\sigma=1$ by a Phragmén–Lindelöf argument (cf. [14, Thm. 8.4]).
Step 1. $Z_{c}/Z_{0}$ is unbounded in $\{\sigma>1,\;|t|\geq 2\}$. For any $N$, by Dirichlet’s approximation theorem there exists $t\leq(2\pi/\varepsilon)^{N}$ with $|\omega_{k}t+\phi_{k}-\pi|<\varepsilon\pmod{2\pi}$ for all $k\leq N$. At $\sigma=1+1/\log N$:
$\operatorname{Re}\log\frac{Z_{c}}{Z_{0}}(\sigma\!+\!it)\geq c\sum_{k=1}^{N}p_{k}^{-\sigma}-O(c)\geq\tfrac{c}{3}\log\log N-O(1),$
which grows without bound as $N\to\infty$ (since $\sum_{p\leq N}p^{-1-1/\log N}\geq e^{-2}\sum_{p\leq N}p^{-1}\sim e^{-2}\log\log N$ by Mertens).
Step 2. Suppose $|Z_{c}/Z_{0}(1+it)|\leq M$ for all $t$. By Phragmén–Lindelöf in the strip $1\leq\sigma\leq 2$ (the function has at most polynomial growth on $\sigma=1$ and is bounded on $\sigma=2$), $Z_{c}/Z_{0}$ would be bounded throughout the strip. This contradicts Step 1.
Therefore $Z_{c}/Z_{0}$ is unbounded on $\sigma=1$. Since $|Z_{0}(1+it)|=O(1)$ (Remark 3.2) and $|Z_{0}(1+it)|$ is bounded away from zero for $|t|\geq 2$, it follows that $\sup_{t}|Z_{c}(1+it)|=+\infty$.
The transition is thus: for $\sigma>1$, $|Z_{c}|\leq\exp(c\,\zeta_{\alpha}(\sigma)+O(1))$ (unconditional by Lemma 1.2); on $\sigma=1$, $\sup_{t}|Z_{c}(1+it)|=+\infty$ (unconditional by the Dirichlet–PLF argument above). (Note: $Z_{c}$ has a pole at $s=1$ from the factor $1/[\tfrac{1}{2}s(s-1)]$; the large values are for $|t|\to\infty$.) ∎
6. The Ratio $Z_{c}/Z_{0}$ and Dirichlet Series
Remark 6.1 (Dirichlet series structure).
Since $\alpha_{k}=p_{k}$, the spectral identity becomes
$\log\frac{Z_{c}}{Z_{0}}(s)=-c\sum_{k=1}^{\infty}e^{-i\phi_{k}}\,p_{k}^{-s}+O(c),$
which is an ordinary Dirichlet series supported on primes. Exponentiating:
(9) $\frac{Z_{c}}{Z_{0}}(s)=\prod_{p}\exp\!\bigl(-c\,e^{-i\phi_{k}}\,p^{-s}\bigr)=\sum_{n=1}^{\infty}a_{n}\,n^{-s},$
with $|a_{p}|=c$,   $|a_{p^{m}}|=c^{m}/m!$,   and $a_{mn}=a_{m}a_{n}$ for $\gcd(m,n)=1$. This converges absolutely for $\sigma>1$. Since both $\Xi_{c}$ and $\Xi_{0}$ satisfy $\Xi(s)=\Xi(1\!-\!s)$ with the same $\Gamma/\pi$ prefactor, the ratio inherits a functional equation:
$\frac{Z_{c}}{Z_{0}}(s)=\frac{Z_{c}}{Z_{0}}(1-s).$
The ratio $Z_{c}/Z_{0}$ is meromorphic: it has zeros at $\{1/2+i\gamma_{n}^{(c)}\}$ (from $Z_{c}$) and poles at $\{1/2+i\gamma_{n}^{(0)}\}$ (from $Z_{0}$), both on $\sigma=1/2$. It is a Dirichlet series (with generalized Euler product) that is bounded for $\sigma>1$ and unbounded on $\sigma=1$, with all its zeros provably on the critical line.
The construction separates the Euler product (which lives on the ratio) from the entireness (which lives on $\Xi_{c}$). Note that this variant uses $\alpha_{k}=p_{k}$ (rational frequencies) and produces a single function $\Xi_{c}$ with the same properties as in Theorem 1.1; the ratio $Z_{c}/Z_{0}$ is then a Dirichlet series, but $Z_{0}$ is not.
7. Almost-Periodicity
The transition at $\sigma=1$ has a natural interpretation in the hierarchy of almost-periodic function spaces.
Proposition 7.1.
Let $\sigma>\tfrac{1}{2}$.
  1. (i)
    For $\sigma>1$: the function $t\mapsto\log|Z_{c}/Z_{0}|(\sigma\!+\!it)$ is Bohr almost-periodic.
  2. (ii)
    For $\sigma>\tfrac{1}{2}$: the function $t\mapsto\log|Z_{c}/Z_{0}|(\sigma\!+\!it)$ is Besicovitch ($B^{2}$) almost-periodic with
    $\|{\log|Z_{c}/Z_{0}|}\|_{B^{2}}^{2}=\tfrac{1}{2}\,c^{2}\sum_{k=1}^{\infty}\alpha_{k}^{-2\sigma}<\infty.$
  3. (iii)
    The construction $E(T)$ is not $B^{2}$-AP: $\sum|a_{k}|^{2}=\frac{1}{\pi^{2}}\sum\alpha_{k}^{-1}=\infty$ (Mertens). The truncation $E_{K}\in B^{2}\text{-AP}$ with $\|E_{K}\|_{B^{2}}^{2}=\frac{1}{2\pi^{2}}\sum_{k=1}^{K}\alpha_{k}^{-1}$.
Proof.
(i) By Proposition 3.1, the function equals an absolutely convergent trigonometric series $c\sum\alpha_{k}^{-\sigma}\cos(\omega_{k}t+\phi_{k})+O(c)$ for $\sigma>1$. Since $\sum\alpha_{k}^{-\sigma}$ converges for $\sigma>1$ [11, Thm. 2.7], the series defines a Bohr-AP function.
(ii) The Fourier–Bohr coefficients at exponents $\pm\omega_{k}$ have squared amplitudes $(c\,\alpha_{k}^{-\sigma}/2)^{2}$. Since $\sum\alpha_{k}^{-2\sigma}$ converges for $2\sigma>1$, the Riesz–Fischer theorem for $B^{2}$-AP functions gives the result.
(iii) The amplitudes of $E(T)$ are $a_{k}=1/(\pi\alpha_{k}^{1/2})$. Since $\sum|a_{k}|^{2}=(1/\pi^{2})\sum\alpha_{k}^{-1}\to\infty$ (by Mertens’ theorem [11, Thm. 2.7]), $E$ does not lie in $B^{2}\text{-AP}$ (its mean square diverges). For the truncation $E_{K}$, the finite sum confirms $E_{K}\in B^{2}\text{-AP}$. ∎
The AP hierarchy for $Z_{c}$ mirrors that of $\zeta$: Bohr-AP for $\sigma>1$ (Bohr [1]), $B^{2}$-AP for $\sigma>1/2$ (Bohr–Jessen [3]), with the passage from convergent to divergent Fourier norm at $\sigma=1$ being the “Bohr-AP to $B^{2}$-AP” transition. The transitions occur at the same critical $\sigma$ values because $\alpha_{k}\sim p_{k}$; the difference is the source: arithmetic (Euler product) for $\zeta$, purely analytic (spectral identity) for $Z_{c}$.
8. Observations
We illustrate the construction numerically with $c=0.05$ and a finite truncation $K=80$. (The computations use the shifted variant $\alpha_{k}=p_{k}+\theta$ with $\theta=e^{\sqrt{2}}$, which ensures global monotonicity of $F_{K}$ (Remark 2.3); the qualitative behavior is identical to the $\alpha_{k}=p_{k}$ case.)
8.1. First zeros of $Z_{c}$
The perturbed zeros $\gamma_{n}$ (defined by $F_{K}(\gamma_{n})=n$ for the truncation) and their shifts $\delta_{n}=\gamma_{n}-\gamma_{n}^{(0)}$:
$n$ $\gamma_{n}$ $\delta_{n}$ gap $E(\gamma_{n})$
1 17.9157 +0.068+0.068 0.011-0.011
2 23.2341 +0.062+0.062 5.318 0.013-0.013
3 27.7063 +0.035+0.035 4.472 0.008-0.008
4 31.6613 0.057-0.057 3.955 +0.015+0.015
5 35.4981 +0.030+0.030 3.837 0.008-0.008
10 51.7580 +0.024+0.024 3.061 0.008-0.008
20 78.6866 +0.014+0.014 2.522 0.006-0.006
50 144.5359 +0.005+0.005 2.038 0.003-0.003
100 236.8981 +0.045+0.045 1.808 0.026-0.026
Max shift: $|\delta_{n}|\leq 0.077$. Min gap: $1.67>0$. Gaps decrease as $\sim 2\pi/\log(\gamma_{n}/(2\pi))$, matching the Weyl law.
8.2. Comparison with $\zeta$ zero statistics
The microscopic statistics of $Z_{c}$’s zeros differ fundamentally from those of $\zeta$. With 1000 zeros computed ($t\in[17.9,1420.1]$):
Statistic $\zeta$ (conditional on RH) $Z_{c}$ Clock
RMS of error term $\sqrt{\frac{\log\log T}{2\pi^{2}}}\to\infty$ $0.013$ (constant) 0
Spacing std (normalized) $\approx 0.42$ (GUE) $0.020$ 0
Pair correlation $1-\mathrm{sinc}^{2}\alpha$ peaks at integers $\delta$’s
Number variance $\Sigma^{2}(L)$ $\sim\tfrac{2}{\pi^{2}}\log L$ $\lesssim 0.25$ 0
Distribution of $S$ / $E$ Gaussian bounded, non-Gaussian 0
The zeros of $Z_{c}$ are super-rigid (nearly crystalline), with normalized spacing std two orders of magnitude below GUE. The construction captures the macroscopic transition (bounded/unbounded, AP hierarchy) but not the microscopic GUE statistics, which arise from the arithmetic content (Euler product, prime correlations) that $Z_{c}$ deliberately lacks.
8.3. Value distribution
Figure 1 shows the complex-plane images of $\log(Z_{c}/Z_{0})(\sigma+it)$, $\log\zeta(\sigma+it)$, and $\log L(\sigma+it,\chi_{4})$ as $t$ ranges over $[20,300]$, for $\sigma=1.25,1.0,0.75$.
The analytic formula gives
(10) $\log\frac{Z_{c}}{Z_{0}}(s)=-c\sum_{k=1}^{\infty}\alpha_{k}^{-s}\,e^{i\phi_{k}}+O(c),$
which has equal real and imaginary RMS: $\langle(\operatorname{Re}\log\frac{Z_{c}}{Z_{0}})^{2}\rangle=\langle(\operatorname{Im}\log\frac{Z_{c}}{Z_{0}})^{2}\rangle=\tfrac{c^{2}}{2}\sum\alpha_{k}^{-2\sigma}$. The value distribution is isotropic in the complex plane, as for $\log\zeta$ via Bohr–Jessen [3].
The quantitative difference is dramatic: at $\sigma=0.75$, the RMS of $\log(Z_{c}/Z_{0})$ is approximately $0.046$, while $\log\zeta$ has RMS $\approx 0.88$. The construction’s value distribution is confined to a bounded disk, whereas $|\log\zeta(\sigma+it)|$ is unbounded. This means $Z_{c}/Z_{0}$ cannot satisfy Voronin universality [15], and its moments do not exhibit the $(\log T)^{k^{2}}$ growth of $\zeta$. The value distribution is isotropic but concentrated—another manifestation of the missing arithmetic.
Figure 1. Complex-plane image of $\log f(\sigma+it)$ for $t\in[20,300]$. Left: $\log(Z_{c}/Z_{0})$; center: $\log\zeta$; right: $\log L(s,\chi_{4})$. Rows: $\sigma=1.25,1.0,0.75$. Color encodes $t$ (yellow==early, purple==late). Note the ${\sim}20{\times}$ difference in scale between the left column and the others.
Figure 2. Shape comparison: $\log(Z_{c}/Z_{0})$ rescaled to match the RMS of $\log\zeta$ (left) vs. $\log\zeta$ (right). The rescaled $Z_{c}$ fills a banded region (reflecting the coherent trigonometric structure) rather than the isotropic spirals of $\log\zeta$ (reflecting multiplicative independence of primes).
8.4. Numerical evidence for the linearization condition
Lemma 1.2 establishes that $\sum_{n=1}^{\infty}E(\gamma_{n}^{(0)})^{2}/n^{2}<\infty$. We verify this numerically with $K=80$ terms and $N=5000$ zero positions.
Growth of $|E(\gamma_{n}^{(0)})|$. The values $|E(\gamma_{n}^{(0)})|$ remain bounded: over the first 5000 zeros, $\max_{n}|E(\gamma_{n}^{(0)})|\approx 0.813$. A power-law regression gives $|E(\gamma_{n}^{(0)})|\sim n^{-0.025}$ (essentially flat), and a logarithmic fit yields $|E(\gamma_{n}^{(0)})|\sim(\log n)^{-0.16}$ (very slightly decreasing). The Abel-summation bound $|E(T)|=O(T)$ grossly overestimates the actual behavior at zero positions.
Mean square and Weyl equidistribution. The empirical mean square agrees with the Weyl prediction:
NN $\frac{1}{N}\sum_{n=1}^{N}E(\gamma_{n}^{(0)})^{2}$ $\frac{1}{2\pi^{2}}\sum_{k=1}^{K}\alpha_{k}^{-1}$ ratio
100 0.0688 0.0654 1.053
500 0.0668 0.0654 1.022
1000 0.0649 0.0654 0.994
5000 0.0648 0.0654 0.991
The ratio stabilizes near $1$, consistent with joint equidistribution of the $K$ phases $\{\omega_{k}\gamma_{n}^{(0)}+\phi_{k}\}$ modulo $2\pi$ (cross terms vanish, diagonal terms give the Mertens sum).
Convergence of $\sum E^{2}/n^{2}$. The partial sums of $\sum_{n=1}^{N}E(\gamma_{n}^{(0)})^{2}/n^{2}$ stabilize rapidly:
NN $\sum_{n=1}^{N}E^{2}/n^{2}$ $\sum_{n=1}^{N}\delta_{n}^{2}/\gamma_{n}^{2}$ last term $E^{2}/n^{2}$
100 0.0891 $3.79\times 10^{-5}$ $2.80\times 10^{-5}$
1000 0.0897 $88\times 10^{-5} 3.3$$30\times 10^{-7}$
5000 0.0898 $3.89\times 10^{-5}$ $3.00\times 10^{-11}$
The growth from $N=1000$ to $N=5000$ is only $0.06\%$. The linearization error $\sum\delta_{n}^{2}/\gamma_{n}^{2}\approx 3.9\times 10^{-5}$ is far below $c^{2}=2.5\times 10^{-3}$.
Spectral identity remainder. The remainder $R(t)$ in the spectral identity $\log|\Xi_{c}/\Xi_{0}|(\sigma\!+\!it)=-c\sum\alpha_{k}^{-\sigma}\cos(\omega_{k}t+\phi_{k})+R(t)$ has:
$\sigma$ $\operatorname{mean}(R)$ $\operatorname{std}(R)$ $\max|R|$
1.50 0.015-0.015 $0.021$ $0.081$
1.10 0.016-0.016 $0.030$ $0.122$
1.01 0.016-0.016 $0.034$ $0.136$
1.00 0.016-0.016 $0.034$ $0.138$
The remainder is $O(c)$ as claimed in Proposition 3.1, and the contribution from linearization ($\sim 10^{-5}$) is negligible compared to the constant and lower-limit terms.
9. Open Problems
Several directions for further investigation suggest themselves.
GUE statistics. The zeros of $Z_{c}$ are super-rigid (nearly crystalline spacing, §8.2), far from the GUE statistics of $\zeta$. What additional structure on the zero perturbation $E$ would force GUE pair correlation onto the zeros of $\Xi_{c}$?
Pointwise bounds on $E$. Numerically, $|E(\gamma_{n}^{(0)})|\leq 1.05$ for all computed nn8.4), vastly smaller than the Abel bound $O(\gamma_{n})$. Proving $|E(\gamma_{n}^{(0)})|=O(1)$ would give a much tighter linearization estimate than Lemma 1.2.
Growth rate on $\sigma=1$. The Phragmén–Lindelöf argument gives $\sup_{t}|Z_{c}(1+it)|=+\infty$ but not the rate. For $\zeta$, $|\zeta(1+it)|=\Omega(\log\log t)$ [14, Thm. 8.5]. Does $|Z_{c}(1+it)|$ achieve the same rate?
10. Concluding Remarks
The function $\Xi_{c}$ provides the first explicit entire function of order $1$ that provably satisfies an analogue of the Riemann Hypothesis (all zeros on $\sigma=1/2$) together with the sharp bounded/unbounded transition at $\sigma=1$ characteristic of $\zeta$. The construction demonstrates that these properties are simultaneously achievable and provable, given sufficient control over the zero distribution.
The ratio $Z_{c}/Z_{0}$ is a Dirichlet series with generalized Euler product, providing a concrete object whose zeros are provably on the critical line and whose analytic properties (boundedness, value distribution, almost-periodicity) can be studied unconditionally. In the context of the Selberg class [12, 9], the construction shows that an RH-like zero distribution is compatible with the $\sigma=1$ transition—the two properties reinforce rather than obstruct each other.
The microscopic statistics of $Z_{c}$’s zeros differ from those of $\zeta$8.2): the zeros are nearly crystalline (super-rigid spacing) rather than GUE-distributed, and $Z_{c}$ does not satisfy Voronin universality. These differences reflect the specific arithmetic content—prime correlations, multiplicative independence of $p^{-it}$—that $\zeta$ possesses and $Z_{c}$ deliberately lacks. Understanding which additional structure would force GUE statistics onto the zeros of an RH-satisfying function remains an interesting open problem.

### v2

@@
 The function thus satisfies an analogue of the Riemann Hypothesis together with the sharp bounded/unbounded transition at $\sigma=1$ characteristic of $\zeta$.
-The construction uses Hadamard products with controlled zero perturbations; the transition is proved unconditionally via a dyadic large-sieve argument and a Phragmén–Lindelöf transfer.
+The transition is controlled by a Dirichlet series $D(s)=\sum e^{-ik\theta}\,p_{k}^{-s}$ whose absolute convergence for $\sigma>1$ and divergence at $\sigma=1$ drive the dichotomy, in direct analogy with $\log\zeta(s)=\sum p^{-s}+O(1)$.
+The key technical input is a dyadic large-sieve estimate establishing the linearization condition that connects the Hadamard product to $D$.
+The construction and proofs were developed in collaboration with Claude (Anthropic); see Acknowledgments.
 1. Introduction
@@
   • For $\sigma>1$: $|\zeta(\sigma+it)|\leq\zeta(\sigma)<\infty$.
-  • On $\sigma=1$: $\limsup_{t\to\infty}|\zeta(1+it)|=\infty$, with $|\zeta(1+it)|=\Omega(\log\log t)$ [14, Thm. 8.5].
+  • On $\sigma=1$: $\limsup_{t\to\infty}|\zeta(1+it)|=\infty$, with $|\zeta(1+it)|=\Omega(\log\log t)$ [17, Thm. 8.5].
 The Selberg class [15, 10] axiomatizes $L$-functions via five properties: Dirichlet series, analytic continuation, functional equation, Ramanujan bound, and Euler product.
@@
 For $\zeta$ itself, the location of zeros is the central open problem; but for a suitably designed function, one can hope to build in the zeros-on-the-line property while retaining enough structure for the $\sigma=1$ transition to hold.
-This paper provides such a construction: an explicit entire function $\Xi_{c}(s)$ of order $1$ with all zeros provably on $\operatorname{Re}(s)=1/2$, satisfying $\Xi_{c}(s)=\Xi_{c}(1-s)$, whose normalized form is uniformly bounded for $\sigma>1$ yet unbounded on $\sigma=1$—the same sharp dichotomy exhibited by $\zeta$.
-The associated ratio $Z_{c}/Z_{0}$ is a Dirichlet series with generalized Euler product, providing a concrete “RH-satisfying” $L$-function–like object whose analytic properties can be studied unconditionally.
+The mechanism behind $\zeta$’s transition is well understood at the level of Dirichlet series: $\log\zeta(s)=P(s)+O(1)$ for $\sigma>1$, where $P(s)=\sum p^{-s}$ is the prime zeta function [17, §1.6, §3.2].
+Boundedness for $\sigma>1$ follows from the absolute convergence of $\sum p^{-\sigma}$; unboundedness on $\sigma=1$ from its divergence (Mertens [14, Thm.
+2.7]).
+The zero-counting error $S(T)=\pi^{-1}\arg\zeta(\tfrac{1}{2}+iT)$ satisfies $S(T)\approx-\pi^{-1}\sum_{p}\sin(T\log p)/\sqrt{p}$ in $L^{2}$ [17, Thm.
+14.22]—the imaginary part of $P(\tfrac{1}{2}+iT)/\pi$.
+Our construction replaces $P(s)$ with a twisted variant $D(s)=\sum e^{-ik\theta}\,p_{k}^{-s}$ (with $\theta\in(0,2\pi)$), uses its restriction to $\sigma=1/2$ to perturb the zeros of a Hadamard product, and proves that the resulting function’s log-ratio with the unperturbed product is approximately cD(s)-cD(s)—the “spectral identity” of this paper.
 Theorem 1.1 (Main Theorem).
-For any $c>0$, define $\Xi_{c}(s)$ as in §2, with $Z_{c}(s)=\Xi_{c}(s)/[\tfrac{1}{2}s(s-1)\pi^{-s/2}\Gamma(s/2)]$. Then:
+Fix $\theta\in(0,2\pi)$ (e.g., $\theta=1$) and $c>0$ sufficiently small. Define $\Xi_{c}(s)$ as in §2, with $Z_{c}(s)=\Xi_{c}(s)/[\tfrac{1}{2}s(s-1)\pi^{-s/2}\Gamma(s/2)]$. Then:
   1. (a)
@@
   4. (d)
-    For each $\delta>0$, $\sup_{t}|Z_{c}(1+\delta+it)|\leq\exp(c\,\zeta_{\alpha}(1+\delta)+O(1))$.
+    For each $\delta>0$, $\sup_{t}|Z_{c}(1+\delta+it)|\leq\exp\bigl(c\,P(1+\delta)+O(1)\bigr)$, where $P(\sigma)=\sum_{p}p^{-\sigma}$.
+Moreover $P(\sigma)\to 0$ as $\sigma\to\infty$, so the bound tends to $e^{O(1)}$: the ratio $\Xi_{c}/\Xi_{0}$ is uniformly close to a constant for $\sigma$ large.
   5. (e)
     $\sup_{t}|Z_{c}(1+it)|=+\infty$.
-Parts (a)–(c) hold by construction; (d) via the spectral identity (Proposition 3.1) with linearization error controlled by Lemma 1.2; and (e) via a Phragmén–Lindelöf argument (§5).
-Moreover, the ratio $Z_{c}/Z_{0}$ is itself a Dirichlet series with generalized Euler product, bounded for $\sigma>1$ and unbounded on $\sigma=1$, with all zeros provably on $\sigma=1/2$.
-Lemma 1.2 (Linearization convergence).
-(1) $\sum_{n=1}^{\infty}\frac{E(\gamma_{n}^{(0)})^{2}}{n^{2}}<\infty,$
-where $E$ is the full (infinite) series (3) evaluated at the regular zero positions.
+  6. (f)
+    The ratio $\Xi_{c}/\Xi_{0}$ admits the approximate Euler product $\Xi_{c}(s)/\Xi_{0}(s)=\kappa_{0}\cdot\prod_{k}\exp(-c\,e^{-ik\theta}p_{k}^{-s})\cdot M(s)$ for $\sigma>1$, where $\kappa_{0}\neq 0$ is an explicit constant, $M(s)$ is bounded (with $M\to 1$ as $c\to 0$), and the Euler product is itself a Dirichlet series with multiplicative coefficients.
+2. The Dirichlet Series and Construction
+2.1. The Dirichlet series $D(s)$
+Fix $\theta\in(0,2\pi)$ (e.g., $\theta=1$), an amplitude $c>0$, and let $p_{1}<p_{2}<\cdots$ denote the primes. Define the Dirichlet series
+(1) $D(s)=\sum_{k=1}^{\infty}e^{-ik\theta}\,p_{k}^{-s},$
+which converges absolutely for $\sigma>1$ (since $\sum p_{k}^{-\sigma}<\infty$) and conditionally for $\sigma>0$, by Abel summation: the partial sums $G_{K}=\sum_{k=1}^{K}e^{-ik\theta}$ satisfy $|G_{K}|\leq 2/|1-e^{-i\theta}|=:C_{\theta}$ (finite since $\theta\in(0,2\pi)$), and the amplitudes $p_{k}^{-\sigma}$ decrease monotonically to zero.
+Viewing $D$ as an ordinary Dirichlet series $\sum b_{n}\,n^{-s}$ (with $b_{n}=e^{-ik\theta}$ for $n=p_{k}$, zero otherwise), the partial sums $B(x)=\sum_{n\leq x}b_{n}$ satisfy $|B(x)|\leq C_{\theta}$, so the abscissa of convergence is $\sigma_{c}=0$ (cf.
+[14, Thm.
+1.3]).
+In the strip of conditional convergence, [14, Thm.
+1.5] gives
+(2) $|D(\sigma+it)|\ll(|t|+4)^{1-\sigma+\varepsilon}$
+for any fixed $\sigma\in(0,1)$ and $\varepsilon>0$. For $\sigma>1$ the stronger bound $|D(\sigma+it)|\leq P(\sigma)=\sum p^{-\sigma}$ holds by absolute convergence.
+Define the perturbation function
+(3) $E(T)=\frac{1}{\pi}\,\operatorname{Im}\bigl[D(\tfrac{1}{2}+iT)\bigr]=-\frac{1}{\pi}\sum_{k=1}^{\infty}\frac{\sin(T\log p_{k}+\phi_{k})}{p_{k}^{1/2}},$
+where $\phi_{k}=k\theta$.
+The convergence of $E(T)$ for each $T>0$ follows from that of $D(1/2+iT)$, and by (2), $|E(T)|=O(|T|^{1/2+\varepsilon})$.
+We write $\omega_{k}=\log p_{k}$ for brevity in Fourier-analytic arguments.
+2.2. The perturbed zeros
+Define the smooth counting function
+(4) $N_{\mathrm{main}}(T)=\frac{T}{2\pi}\log\frac{T}{2\pi}-\frac{T}{2\pi}+\frac{7}{8},\qquad N_{\mathrm{main}}^{\prime}(T)=\frac{1}{2\pi}\log\frac{T}{2\pi}.$
+Let $\{\gamma_{n}^{(0)}\}$ be the “regular” sequence defined by $N_{\mathrm{main}}(\gamma_{n}^{(0)})=n$, with Hadamard product $\Xi_{0}(s)$.
+Define the “perturbed” sequence $\gamma_{n}=\gamma_{n}^{(0)}+\delta_{n}$ with
+(5) $\delta_{n}=-\frac{cE(\gamma_{n}^{(0)})}{N_{\mathrm{main}}^{\prime}(\gamma_{n}^{(0)})},$
+and let $\Xi_{c}(s)$ be the Hadamard product (cf. [6, §1.10])
+(6) $\Xi_{c}(s)=\prod_{n=1}^{\infty}\left(1+\frac{(s-\tfrac{1}{2})^{2}}{\gamma_{n}^{2}}\right).$
+Lemma 2.1 (Well-definedness).
+For $c$ sufficiently small, $\gamma_{n}>0$ for all $n$, $\sum\gamma_{n}^{-2}<\infty$, and $\Xi_{c}(s)$ is an entire function of order $1$.
+Proof.
+Since $|E(T)|=O(|T|^{1/2+\varepsilon})$, $|\delta_{n}|=O(c(\gamma_{n}^{(0)})^{1/2+\varepsilon}/\log\gamma_{n}^{(0)})=o(\gamma_{n}^{(0)})$, so $\gamma_{n}>\gamma_{n}^{(0)}/2$ for all $n$ sufficiently large.
+For the finitely many small $n$, $\delta_{n}$ is finite (since $E(\gamma_{n}^{(0)})$ converges for each $n$) and proportional to $c$, so taking $c$ small enough ensures $\gamma_{n}>0$ for all $n$.
+Since $\gamma_{n}^{(0)}\sim 2\pi n/\log n$, $\sum\gamma_{n}^{-2}<\infty$ and the exponent of convergence equals $1$.
+∎
+We set $Z_{c}(s)=\Xi_{c}(s)/\Gamma_{\pi}(s)$ and $Z_{0}(s)=\Xi_{0}(s)/\Gamma_{\pi}(s)$, where $\Gamma_{\pi}(s)=\tfrac{1}{2}s(s-1)\pi^{-s/2}\Gamma(s/2)$.
+Note that $Z_{c}/Z_{0}=\Xi_{c}/\Xi_{0}$ (the $\Gamma_{\pi}$ factors cancel).
+3. The Spectral Identity
+The central result relates the Hadamard product ratio $\Xi_{c}/\Xi_{0}$ to the Dirichlet series $D$.
+Proposition 3.1 (Spectral Identity).
+Let $\omega_{k}=\log p_{k}$ and $\phi_{k}=k\theta$ as in §2. For $\sigma>1/2$ and $|t|\geq 2$:
+(7) $\operatorname{Re}\log\frac{\Xi_{c}}{\Xi_{0}}(\sigma+it)=-c\sum_{k}p_{k}^{-\sigma}\cos(\omega_{k}t+\phi_{k})+O(1).$
+Equivalently, there exists a bounded analytic function $R(s)$ with $R(s)=O(1)$ uniformly for $\sigma>1/2$, $|t|\geq 2$, such that
+(8) $\log\frac{\Xi_{c}}{\Xi_{0}}(s)=-c\,D(s)+R(s).$
+The cosine sum converges absolutely for $\sigma>1$ and conditionally for $\sigma>0$ (as $\operatorname{Re}\,D(\sigma+it)$).
+Proof.
+We establish (7) first, then deduce (8) by analyticity. Throughout, $c$ is fixed, so all implicit constants may depend on $c$ and $\theta$.
+Step 1 (Linearization). By (5), $\gamma_{n}=\gamma_{n}^{(0)}+\delta_{n}$. Each Hadamard factor satisfies:
+$\log\frac{1+(s-\tfrac{1}{2})^{2}/\gamma_{n}^{2}}{1+(s-\tfrac{1}{2})^{2}/(\gamma_{n}^{(0)})^{2}}=\frac{-2w^{2}\delta_{n}}{\gamma_{n}^{(0)}((\gamma_{n}^{(0)})^{2}+w^{2})}+O(\delta_{n}^{2}/(\gamma_{n}^{(0)})^{2})$
+where w=s1/2w=s-1/2.
+Step 2 (Sum to integral).
+For each $K$, let $\Xi_{c,K}$ be the Hadamard product with zeros perturbed using $E_{K}$ (first $K$ terms of $E$).
+Since $|E_{K}|\leq\frac{1}{\pi}\sum_{k=1}^{K}p_{k}^{-1/2}$ (a constant for each $K$), the perturbations are $\delta_{n,K}=O(1/\log\gamma_{n}^{(0)})$, giving convergent quadratic remainder $\sum\delta_{n,K}^{2}/(\gamma_{n}^{(0)})^{2}<\infty$.
+The linearized sum is a Riemann sum for $\int\frac{2w^{2}\,cE_{K}(u)}{u(u^{2}+w^{2})}\,du$ (one term per zero, spacing $\sim 1/N_{\mathrm{main}}^{\prime}$).
+The integrand is smooth for $u>0$ and $\sigma>1/2$ (the poles of $1/(u^{2}+w^{2})$ lie off the real axis), so by Euler–Maclaurin:
+(9) $\operatorname{Re}\log\frac{\Xi_{c,K}(s)}{\Xi_{0}(s)}=-c\int_{\gamma_{1}^{(0)}}^{\infty}\operatorname{Re}\frac{2w^{2}\,E_{K}(u)}{u(u^{2}+w^{2})}\,du+O(1).$
+Step 3 (Evaluation of the integral).
+Using the partial fraction $\frac{2w^{2}}{u(u^{2}+w^{2})}=\frac{2}{u}-\frac{2u}{u^{2}+w^{2}}$ and the standard integrals $\int_{0}^{\infty}\frac{\sin\omega u}{u}\,du=\frac{\pi}{2}$ (Dirichlet) and $\int_{0}^{\infty}\frac{u\sin\omega u}{u^{2}+\beta^{2}}\,du=\frac{\pi}{2}e^{-\omega\beta}$ for $\omega>0$, $\operatorname{Re}\beta>0$ ([8, §3.723], extended to complex $\beta$ by analytic continuation), the contribution of each component $\sin(\omega_{k}u+\phi_{k})$ decomposes into convergent pieces.
+[^1] The individual integrals $\int\sin(\omega u+\phi)/u\,du$ and $\int u\sin(\omega u+\phi)/(u^{2}+w^{2})\,du$ each diverge at $u=0$ when $\sin\phi\neq 0$ (from the $\cos(\omega u)/u$ term), but the divergent parts cancel in the partial-fraction difference.
+The oscillatory contribution of the $k$-th component is $\pi\,p_{k}^{-(\sigma-1/2)}\cos(\omega_{k}t+\phi_{k})$ (using $e^{-\omega_{k}(\sigma-1/2)}=p_{k}^{-(\sigma-1/2)}$ and the angle-addition identity); the non-oscillatory part $\pi\cos\phi_{k}(1-p_{k}^{-(\sigma-1/2)})$ sums to a bounded quantity (convergent by Abel summation, depending on $\sigma$ but not on $t$).
+Substituting into (9) with the amplitudes $1/(\pi p_{k}^{1/2})$ from $E_{K}$, the $k$-th term contributes $c\,p_{k}^{-\sigma}\cos(\omega_{k}t+\phi_{k})$.
+Including the bounded lower-limit correction $\int_{0}^{\gamma_{1}^{(0)}}(\cdots)\,du$:
+(10) $\operatorname{Re}\log\frac{\Xi_{c,K}(s)}{\Xi_{0}(s)}=-c\sum_{k=1}^{K}p_{k}^{-\sigma}\cos(\omega_{k}t+\phi_{k})+O(1)+O\Bigl(\sum_{n}\delta_{n,K}^{2}/(\gamma_{n}^{(0)})^{2}\Bigr),$
+where the $O(1)$ accounts for the non-oscillatory Fourier terms, the lower-limit correction, and the Euler–Maclaurin error (all independent of $K$), while the final term is the quadratic Taylor remainder from Step 1.
+For each fixed $K$ this remainder is finite, but it grows with $K$ (since $M_{K}$ grows).
+Taking $K\to\infty$: the cosine sum converges (absolutely for $\sigma>1$; conditionally for $\sigma>0$), and $\Xi_{c,K}\to\Xi_{c}$ (since the zero sequences converge).
+The only obstruction to passing (10) to the limit is the Taylor remainder: for the full (non-truncated) series, $\delta_{n}=-cE(\gamma_{n}^{(0)})/N_{\mathrm{main}}^{\prime}(\gamma_{n}^{(0)})$ and $|E(\gamma_{n}^{(0)})|$ is no longer bounded by a fixed constant (its RMS grows like $\sqrt{\log\log n}$ by (14)).
+The quadratic remainder $\sum\delta_{n}^{2}/(\gamma_{n}^{(0)})^{2}$ converges iff $\sum E(\gamma_{n}^{(0)})^{2}/n^{2}<\infty$, which is the content of Lemma 3.2 below.
+This establishes (7).
+Step 4 (Complex extension).
+Equation (7) controls the real part of $\log(\Xi_{c}/\Xi_{0})+cD$; we now upgrade to the full complex identity.
+Define $F(s)=\log(\Xi_{c}(s)/\Xi_{0}(s))+cD(s)$.
+Both terms are analytic for $\sigma>1/2$: the Hadamard ratio $\Xi_{c}/\Xi_{0}$ is nonvanishing there (all zeros of $\Xi_{0}$ and $\Xi_{c}$ lie on $\sigma=1/2$), and $D$ converges for $\sigma>0$.
+By (7) (proved above without reference to (8)), $\operatorname{Re}F(s)=O(1)$ for $\sigma>1/2$, $|t|\geq 2$.
+Evaluating at $s=2$: $D(2)$ converges absolutely, and $\log(\Xi_{c}(2)/\Xi_{0}(2))$ is finite (each Hadamard factor at $s=2$ differs from $1$ by $O(\delta_{n}/\gamma_{n}^{2})$, and $\sum|\delta_{n}|/\gamma_{n}^{2}<\infty$), so $F(2)=O(1)$.
+Since $F$ is analytic with bounded real part, the Borel–Carathéodory theorem gives $|F(s)|=O(1)$ for $\sigma>1/2$.
+Setting $R(s)=F(s)$ gives (8).
+∎
+3.1. Linearization convergence
+The following lemma ensures that the Taylor linearization in Step 1 has convergent remainder for the full (non-truncated) series, which is the key input for passing the spectral identity from finite truncations to $\Xi_{c}$ itself.
+Lemma 3.2.
+(11) $\sum_{n=1}^{\infty}\frac{E(\gamma_{n}^{(0)})^{2}}{n^{2}}<\infty.$
 Proof.
@@
 $E_{K_{j}}$ is a trigonometric polynomial with $K_{j}$ terms and frequency range $\Lambda_{j}=\omega_{K_{j}}-\omega_{1}=O(j)$.
-The zeros $\gamma_{n}^{(0)}$ have spacing $\delta\geq c_{1}/j$ for $n\leq 2^{j+1}$.
+The zeros $\gamma_{n}^{(0)}$ have minimum spacing $\Delta\geq c_{1}/j$ for $n\leq 2^{j+1}$.
 By the large sieve inequality for trigonometric polynomials at well-spaced points (cf.
-[10, Thm.
-1], [11, §7]):
-$\sum_{n\leq 2^{j+1}}\!|E_{K_{j}}(\gamma_{n}^{(0)})|^{2}\leq(\pi/\delta+\Lambda_{j})\,\tfrac{1}{\pi^{2}}\sum_{k\leq K_{j}}\alpha_{k}^{-1}\leq C_{1}\,j\log j,$
-since $\sum_{k\leq K_{j}}\alpha_{k}^{-1}\sim\log\log p_{K_{j}}\sim\log j$ (Mertens). Therefore $\text{HEAD}_{j}\leq C_{1}\,j\log j\,/\,4^{j}$, and $\sum_{j}j\log j/4^{j}<\infty$.
+[13, Thm.
+1], [14, §7]):
+$\sum_{n\leq 2^{j+1}}\!|E_{K_{j}}(\gamma_{n}^{(0)})|^{2}\leq(\pi/\Delta+\Lambda_{j})\,\tfrac{1}{\pi^{2}}\sum_{k\leq K_{j}}p_{k}^{-1}\leq C_{1}\,j\log j,$
+since $\sum_{k\leq K_{j}}p_{k}^{-1}\sim\log\log p_{K_{j}}\sim\log j$ (Mertens). Therefore $\text{HEAD}_{j}\leq C_{1}\,j\log j\,/\,4^{j}$, and $\sum_{j}j\log j/4^{j}<\infty$.
 Tail (Abel summation).
-From the proof of Lemma 2.1, the tail $D_{K}=\sum_{k>K}(\omega_{k+1}-\omega_{k})/\alpha_{k}^{1/2}$ is a left Riemann sum for the decreasing function $e^{-\omega/2}$.
-By Bertrand’s postulate (mesh width $\leq\log 2$): $D_{K}\leq\sqrt{2}\int_{\omega_{K}}^{\infty}e^{-\omega/2}\,d\omega=2\sqrt{2}\,\alpha_{K}^{-1/2}\leq 2\sqrt{2}/\sqrt{p_{K}}$.
-Hence $|R_{K_{j}}(\gamma_{n}^{(0)})|\leq(C_{\theta}/\pi)(\alpha_{K_{j}+1}^{-1/2}+\gamma_{n}D_{K_{j}})\leq C_{2}\,2^{j(1-\varepsilon)/2}/j^{3/2}$ (with $\varepsilon=0.1$, using $\gamma_{n}\leq C_{3}\,2^{j}/j$ and $\sqrt{p_{K_{j}}}\geq C_{4}\,2^{(1+\varepsilon)j/2}\sqrt{j}$).
+The tail $\sum_{k>K}(\omega_{k+1}-\omega_{k})/p_{k}^{1/2}$ is a left Riemann sum for the decreasing function $e^{-\omega/2}$.
+By Bertrand’s postulate (mesh width $\leq\log 2$): $\sum_{k>K}(\omega_{k+1}-\omega_{k})/p_{k}^{1/2}\leq\sqrt{2}\int_{\omega_{K}}^{\infty}e^{-\omega/2}\,d\omega=2\sqrt{2}\,p_{K}^{-1/2}$.
+Hence $|R_{K_{j}}(\gamma_{n}^{(0)})|\leq C_{2}\,2^{j(1-\varepsilon)/2}/j^{3/2}$ (with $\varepsilon=0.1$, using $\gamma_{n}\leq C_{3}\,2^{j}/j$ and $\sqrt{p_{K_{j}}}\geq C_{4}\,2^{(1+\varepsilon)j/2}\sqrt{j}$).
 Therefore $\text{TAIL}_{j}\leq C_{5}/(j^{3}\cdot 2^{\varepsilon j})$, and $\sum_{j}1/(j^{3}\cdot 2^{\varepsilon j})<\infty$.
 Both contributions decay exponentially, so $\sum S_{j}<\infty$. ∎
-1.1. Method
-Given a sequence of positive reals $0<t_{1},t_{2},\ldots$ with $\sum t_{n}^{-2}<\infty$, define the Hadamard product
-(2) $\Xi(s)=\prod_{n=1}^{\infty}\left(1+\frac{(s-\tfrac{1}{2})^{2}}{t_{n}^{2}}\right).$
-This has zeros at $s=1/2\pm it_{n}$ and satisfies $\Xi(s)=\Xi(1-s)$ (cf.
-[5, §1.10], [14, §2.12]).
-We choose the zero positions $\{t_{n}\}$ so that the counting function $N_{c}(T)=\#\{n:t_{n}\leq T\}$ approximates $N_{\mathrm{main}}(T)+cE(T)$, where $E$ is a trigonometric series whose Fourier amplitudes decay like the prime reciprocals.
-The growth of $|Z|$ on vertical lines is controlled by the Poisson extension of $E$: bounded for width $w>1/2$ (i.e., $\sigma>1$) and unbounded at width $w=1/2$ (i.e., $\sigma=1$), from a marginal convergence/divergence condition on the damped Fourier amplitudes.
-2. The Explicit Construction
-2.1. Parameters
-Fix an irrational $\theta$ (e.g., $\theta=e^{\sqrt{2}}$), an amplitude $c>0$, and the primes $p_{1}<p_{2}<\cdots$. Define:
-(3) $\alpha_{k}=p_{k},\quad\omega_{k}=\log p_{k},\quad\phi_{k}=k\theta,\quad E(T)=-\frac{1}{\pi}\sum_{k=1}^{\infty}\frac{\sin(\omega_{k}T+\phi_{k})}{p_{k}^{1/2}}.$
-Lemma 2.1 (Convergence).
-The series $E(T)$ converges for all $T>0$.
+3.2. The base function $Z_{0}$
+Lemma 3.3.
+$\log|Z_{0}(\sigma+it)|=O(1)$ uniformly for $\sigma\geq 1$ and $|t|\geq 2$. In particular, $|Z_{0}|$ is both bounded and bounded away from zero in this region.
 Proof.
-Write $E(T)=-(1/\pi)\operatorname{Im}\bigl[\sum_{k}d_{k}\,c_{k}\bigr]$ where $c_{k}=e^{ik\theta}$ and $d_{k}=e^{i\omega_{k}T}/\alpha_{k}^{1/2}$.
-Apply summation by parts (Abel’s formula): $\sum_{k=1}^{K}c_{k}d_{k}=G_{K}d_{K}-\sum_{k=1}^{K-1}G_{k}(d_{k+1}-d_{k})$, where $G_{k}=\sum_{j=1}^{k}e^{ij\theta}$ satisfies $|G_{k}|\leq 2/|1-e^{i\theta}|=:C_{\theta}$ (since $\theta$ is irrational).
-The boundary term satisfies $|G_{K}d_{K}|\leq C_{\theta}\,\alpha_{K}^{-1/2}\to 0$.
-For the variation sum: $|d_{k+1}-d_{k}|\leq|\omega_{k+1}-\omega_{k}|\,T/\alpha_{k}^{1/2}+(\alpha_{k}^{-1/2}-\alpha_{k+1}^{-1/2})$.
-The amplitude differences telescope: $\sum(\alpha_{k}^{-1/2}-\alpha_{k+1}^{-1/2})=\alpha_{1}^{-1/2}$.
-The phase differences satisfy $\sum_{k}(\omega_{k+1}-\omega_{k})/\alpha_{k}^{1/2}=\sum_{k}\log(\alpha_{k+1}/\alpha_{k})/\alpha_{k}^{1/2}<\infty$ (this sum is a Riemann sum for $\int_{\omega_{1}}^{\infty}e^{-\omega/2}\,d\omega<\infty$, with bounded mesh by Bertrand’s postulate).
-Hence $\sum|G_{k}(d_{k+1}-d_{k})|$ converges absolutely, and $G_{K}d_{K}\to 0$, so $\sum c_{k}d_{k}$ converges.
+By construction, $\Xi_{0}$ has zero-counting function $N_{0}(T)=N_{\mathrm{main}}(T)$ (no error term).
+By Euler–Maclaurin, $\log|\Xi_{0}(s)|=\int_{\gamma_{1}^{(0)}}^{\infty}\log|1+w^{2}/u^{2}|\,N_{\mathrm{main}}^{\prime}(u)\,du+O(1)$.
+Stirling’s formula identifies this integral with $\log|\Gamma_{\pi}(s)|+O(1)$, giving $\log|Z_{0}|=\log|\Xi_{0}/\Gamma_{\pi}|=O(1)$.
+The second claim follows since $|\!\log|Z_{0}|\!|\leq C$ implies $e^{-C}\leq|Z_{0}|\leq e^{C}$.
 ∎
-2.2. The perturbed zeros
-Define the smooth counting function
-(4) $N_{\mathrm{main}}(T)=\frac{T}{2\pi}\log\frac{T}{2\pi}-\frac{T}{2\pi}+\frac{7}{8},\qquad N_{\mathrm{main}}^{\prime}(T)=\frac{1}{2\pi}\log\frac{T}{2\pi}.$
-Let $\{\gamma_{n}^{(0)}\}$ be the “regular” sequence defined by $N_{\mathrm{main}}(\gamma_{n}^{(0)})=n$, with Hadamard product $\Xi_{0}(s)$.
-Define the “perturbed” sequence $\gamma_{n}=\gamma_{n}^{(0)}+\delta_{n}$ with
-(5) $\delta_{n}=-\frac{cE(\gamma_{n}^{(0)})}{N_{\mathrm{main}}^{\prime}(\gamma_{n}^{(0)})},$
-and let $\Xi_{c}(s)$ be the Hadamard product (2) with $t_{n}=\gamma_{n}$.
-Lemma 2.2 (Well-definedness of the Hadamard product).
-For any $c>0$, the sequence $\{\gamma_{n}\}$ satisfies $\gamma_{n}>0$ for all $n$ (taking $c$ small enough for finitely many small $n$) and $\sum\gamma_{n}^{-2}<\infty$, so $\Xi_{c}(s)$ is an entire function of order $1$.
-Proof.
-Since $E(T)$ converges for each $T$ (Lemma 2.1), $\delta_{n}$ is finite for each $n$.
-For $n$ large, $|\delta_{n}|=O(c|E(\gamma_{n}^{(0)})|/\log\gamma_{n}^{(0)})$.
-By the Abel-summation bound of Lemma 2.1, $|E(T)|=O(T)$ for each $T$ (the bound grows at most linearly since the total variation of $e^{i\omega_{k}T}$ over $k\leq K$ is $O(T\log\alpha_{K})$, controlled by the monotone decay of $\alpha_{k}^{-1/2}$).
-Thus $|\delta_{n}|=O(c\gamma_{n}^{(0)}/\log\gamma_{n}^{(0)})<\gamma_{n}^{(0)}/2$ for $n$ sufficiently large, giving $\gamma_{n}>\gamma_{n}^{(0)}/2>0$.
-Since $\gamma_{n}^{(0)}\sim 2\pi n/\log n$, we have $\sum\gamma_{n}^{-2}\leq 4\sum(\gamma_{n}^{(0)})^{-2}<\infty$.
-The exponent of convergence equals $1$.
-∎
-Remark 2.3.
-The zeros $\gamma_{n}$ need not be in increasing order: crossings ($\gamma_{n+1}<\gamma_{n}$) may occur where $|E^{\prime}|$ is large.
-For a finite truncation $E_{K}$, global monotonicity of $F_{K}(T)=N_{\mathrm{main}}(T)+cE_{K}(T)$ holds for $c<\log(T_{0}/(2\pi))/(2\pi C_{K})$ with $C_{K}=\frac{1}{\pi}\sum_{k=1}^{K}\omega_{k}\alpha_{k}^{-1/2}$.
-For the infinite series, occasional crossings produce zeros of multiplicity >1>1 without affecting the Hadamard product convergence or the spectral identity (which depends only on the linearization at $\gamma_{n}^{(0)}$, not on the ordering of $\gamma_{n}$).
-3. The Spectral Identity
-The following proposition relates the Hadamard product ratio $\Xi_{c}/\Xi_{0}$ to the Fourier content of $E$. All quantities ($\alpha_{k},\omega_{k},\phi_{k},E$) are as defined in §2.
-Proposition 3.1 (Spectral Identity).
-For $\sigma\geq 1$ and $|t|\geq 2$:
-(6) $\log\frac{|\Xi_{c}(\sigma+it)|}{|\Xi_{0}(\sigma+it)|}=-c\sum_{k}\alpha_{k}^{-\sigma}\cos(\omega_{k}t+\phi_{k})+O(c).$
-Proof.
-Step 1 (Linearization). By (5), $\gamma_{n}=\gamma_{n}^{(0)}+\delta_{n}$. Each Hadamard factor satisfies (Taylor-expanding in $\delta_{n}$):
-$\log\frac{1+(s-\tfrac{1}{2})^{2}/\gamma_{n}^{2}}{1+(s-\tfrac{1}{2})^{2}/(\gamma_{n}^{(0)})^{2}}=\frac{-2w^{2}\delta_{n}}{\gamma_{n}^{(0)}((\gamma_{n}^{(0)})^{2}+w^{2})}+O(\delta_{n}^{2}/(\gamma_{n}^{(0)})^{2})$
-where w=s1/2w=s-1/2.
-Step 2 (Finite-truncation limit).
-For each $K$, let $\Xi_{c,K}$ be the Hadamard product with zeros defined using the truncation $E_{K}$ (first $K$ terms of $E$).
-The perturbations $\delta_{n,K}=-cE_{K}(\gamma_{n}^{(0)})/N_{\mathrm{main}}^{\prime}(\gamma_{n}^{(0)})$ satisfy $|\delta_{n,K}|\leq 2\pi cM_{K}/\log\gamma_{n}^{(0)}$ where $M_{K}=\frac{1}{\pi}\sum_{k=1}^{K}\alpha_{k}^{-1/2}$ is a constant.
-The linearization error for the $K$-truncation is then
-(7) $\log\frac{\Xi_{c,K}(s)}{\Xi_{0}(s)}=-c\int_{\gamma_{1}^{(0)}}^{\infty}\frac{2w^{2}\,E_{K}(u)}{u(u^{2}+w^{2})}\,du+O(c^{2}),$
-where $O(c^{2})$ accounts for the quadratic remainder $\sum\delta_{n,K}^{2}/(\gamma_{n}^{(0)})^{2}$, which converges since $|\delta_{n,K}|=O(cM_{K}/\log\gamma_{n}^{(0)})$ and $\sum 1/(\gamma_{n}^{(0)}\log\gamma_{n}^{(0)})^{2}<\infty$.
-This error is $O(c^{2}M_{K}^{2})$, independent of $t$.
-Since $E_{K}\to E$ pointwise and $\gamma_{n,K}\to\gamma_{n}$ for each $n$, the Hadamard products converge: $\Xi_{c,K}(s)\to\Xi_{c}(s)$ for each $s$, by dominated convergence of the product factors.
-Taking $K\to\infty$ in (7):
-Step 3 (Contour integration).
-For a single component $E_{k}(u)=-\frac{1}{\pi}\alpha_{k}^{-1/2}\sin(\omega_{k}u+\phi_{k})$, we evaluate $I_{k}=\int_{0}^{\infty}\frac{2w^{2}\sin(\omega_{k}u+\phi_{k})}{u(u^{2}+w^{2})}\,du$ by contour integration.
-(The lower limit in (7) is $\gamma_{1}^{(0)}\approx 17.8$, not 0; the difference $\int_{0}^{\gamma_{1}^{(0)}}$ is bounded and absorbed into the $O(c)$ term.)
-Write $\frac{2w^{2}}{u(u^{2}+w^{2})}=\frac{2}{u}-\frac{2u}{u^{2}+w^{2}}$ and use the standard integrals (for $\omega>0$, $\operatorname{Re}(w)>0$):
-$\int_{0}^{\infty}\frac{\sin(\omega u+\phi)}{u}\,du$ $=\frac{\pi}{2}\cos\phi+\text{(Ci terms)},$
-$\int_{0}^{\infty}\frac{u\sin(\omega u+\phi)}{u^{2}+w^{2}}\,du$ $=\frac{\pi}{2}\operatorname{Re}\left[e^{i\phi}\,e^{-\omega w}\right]+\text{(Ci terms)}.$
-The Ci (cosine integral) terms cancel in the difference. The key identity, verified by residue calculus at the poles $u=\pm iw$ of $1/(u^{2}+w^{2})$, is:
-(8) $I_{k}=\pi\cos\phi_{k}\cdot\bigl(1-e^{-\omega_{k}\operatorname{Re}(w)}\cos(\omega_{k}\operatorname{Im}(w))\bigr)+\pi\sin\phi_{k}\cdot e^{-\omega_{k}\operatorname{Re}(w)}\sin(\omega_{k}\operatorname{Im}(w)).$
-For $s=\sigma+it$: $\operatorname{Re}(w)=\sigma-1/2$, $\operatorname{Im}(w)=t$.
-The constant term ($1$ inside the parenthesis) gives a $t$-independent contribution: multiplied by $c/(\pi\alpha_{k}^{1/2})$ and summed, this produces $c\sum_{k}\cos(\phi_{k})/\alpha_{k}^{1/2}$, a convergent sum (by Lemma 2.1 with $T=0$) of magnitude $O(c)$.
-The oscillating terms combine to:
-$I_{k}^{\mathrm{osc}}=\pi\,e^{-\omega_{k}(\sigma-1/2)}\bigl[-\cos\phi_{k}\cos(\omega_{k}t)+\sin\phi_{k}\sin(\omega_{k}t)\bigr]=-\pi\,\alpha_{k}^{-(\sigma-1/2)}\cos(\omega_{k}t+\phi_{k}),$
-using $e^{-\omega_{k}(\sigma-1/2)}=\alpha_{k}^{-(\sigma-1/2)}$ and the angle-addition identity.
-Substituting into (7) with $E_{K}=-\frac{1}{\pi}\sum_{k=1}^{K}\alpha_{k}^{-1/2}\sin(\omega_{k}u+\phi_{k})$, the contribution of the $k$-th term is $c\,\alpha_{k}^{-\sigma}\cos(\omega_{k}t+\phi_{k})$ (the factors of $\pi$ cancel).
-Summing over $k=1,\ldots,K$:
-$\operatorname{Re}\log\frac{\Xi_{c,K}(s)}{\Xi_{0}(s)}=-c\sum_{k=1}^{K}\alpha_{k}^{-\sigma}\cos(\omega_{k}t+\phi_{k})+O(c).$
-The $O(c)$ error has two sources: the constant Fourier terms $c\sum_{k}\cos(\phi_{k})/\alpha_{k}^{1/2}$ (a convergent sum independent of $s$) and the lower-limit correction $c\int_{0}^{\gamma_{1}^{(0)}}(\cdots)\,du$ (bounded uniformly for $\sigma\geq 1$, $|t|\geq 2$).
-Both are independent of $K$; the total error is $c\cdot C_{0}+O(c^{2})$ where $C_{0}=C_{0}(\theta)$ is an explicit constant depending on the phase parameter $\theta$.
-Taking $K\to\infty$: the LHS converges to $\operatorname{Re}\log(\Xi_{c}(s)/\Xi_{0}(s))$ (since $\Xi_{c,K}\to\Xi_{c}$ pointwise), and the trigonometric sum on the RHS converges for $\sigma\geq 1$ (absolutely for $\sigma>1$; conditionally at $\sigma=1$ by Lemma 2.1).
-The linearization error $\sum\delta_{n}^{2}/(\gamma_{n}^{(0)})^{2}$ for the full series converges by Lemma 1.2, contributing $O(c^{2})$.
-The identity (6) follows with total error $O(c)$.
-∎
-Lemma 3.2.
-$\log|Z_{0}(\sigma+it)|=O(1)$ uniformly for $\sigma\geq 1$ and $|t|\geq 2$.
-Proof.
-Since $N_{0}(T)=N_{\mathrm{main}}(T)$ (no error term), the Euler–Maclaurin formula converts the zero sum to an integral: $\log|\Xi_{0}(s)|=\sum_{n}\log|1+w^{2}/(\gamma_{n}^{(0)})^{2}|=\int_{\gamma_{1}^{(0)}}^{\infty}\log|1+w^{2}/u^{2}|\,N_{\mathrm{main}}^{\prime}(u)\,du+O(1)$.
-By Stirling’s formula, $\int\log|1+w^{2}/u^{2}|\,N_{\mathrm{main}}^{\prime}(u)\,du=\log|\Gamma_{\pi}(s)|+O(1)$ where $\Gamma_{\pi}(s)=\tfrac{1}{2}s(s-1)\pi^{-s/2}\Gamma(s/2)$ (the integral reproduces the Weierstrass product of $\Gamma$ at the Weyl density).
-Hence $\log|Z_{0}|=\log|\Xi_{0}/\Gamma_{\pi}|=O(1)$.
-Note that $Z_{0}$ has a simple pole at $s=1$ from $1/[\tfrac{1}{2}s(s-1)]$; the bound is for $|t|\geq 2$.
-∎
-Remark 3.3.
-Consequently, $\log|Z_{c}(\sigma+it)|=-c\sum\alpha_{k}^{-\sigma}\cos(\omega_{k}t+\phi_{k})+O(1)$ for $\sigma\geq 1$, $|t|\geq 2$.
-4. The Fourier-Analytic Mechanism
-The bounded/unbounded transition has a simple Fourier-analytic explanation.
-The spectral identity (Proposition 3.1) gives $\log|Z_{c}/Z_{0}|(\sigma+it)=-c\sum\alpha_{k}^{-\sigma}\cos(\omega_{k}t+\phi_{k})+O(c)$.
-For $\sigma>1$: the sum $\sum\alpha_{k}^{-\sigma}$ converges (since $\alpha_{k}=p_{k}$ and $\sum p^{-\sigma}<\infty$), so the triangle inequality gives $|\log|Z_{c}/Z_{0}||\leq c\sum\alpha_{k}^{-\sigma}+O(c)$—boundedness.
-At $\sigma=1$: the sum $\sum\alpha_{k}^{-1}$ diverges (Mertens [11, Thm.
-2.7]), and the Dirichlet series structure of $Z_{c}/Z_{0}$ (Remark 6.1) enables a Phragmén–Lindelöf argument (§5) to transfer the divergence into unboundedness of $Z_{c}$ on $\sigma=1$.
-The transition is thus driven entirely by the marginal convergence/divergence of the damped Fourier amplitudes $\sum\alpha_{k}^{-\sigma}$ at $\sigma=1$.
 4. Proof of the Main Theorem
 Proof of Theorem 1.1.
-The function $\Xi_{c}$ is the Hadamard product (2) with zeros $\gamma_{n}=\gamma_{n}^{(0)}+\delta_{n}$ defined via the full series $E$ as in (5).
-By Lemma 2.2, $\Xi_{c}$ is an entire function of order $1$.
+Parts (a)–(c) hold by construction; (d) follows from the spectral identity (Proposition 3.1) with linearization error controlled by Lemma 3.2; and (e) via a Phragmén–Lindelöf argument below.
+The function $\Xi_{c}$ is the Hadamard product (6) with zeros $\gamma_{n}=\gamma_{n}^{(0)}+\delta_{n}$ defined via $E$ as in (5). By Lemma 2.1, $\Xi_{c}$ is an entire function of order $1$.
 Part (a): by construction, zeros lie at $s=1/2\pm i\gamma_{n}$. Part (b): $\Xi_{c}(1-s)=\Xi_{c}(s)$ (since $(1-s-1/2)^{2}=(s-1/2)^{2}$).
-Part (c): the counting function $N_{c}(T)=\#\{n:\gamma_{n}\leq T\}$ satisfies $|N_{c}(T)-N_{\mathrm{main}}(T)|=O(\log T)$ because the shifts $\delta_{n}=O(c\gamma_{n}^{(0)}/\log\gamma_{n}^{(0)})$ (Lemma 2.2) can change $N_{c}(T)$ by at most $O(\log T)$ relative to $N_{\mathrm{main}}(T)$.
-Part (d) (Boundedness for $\sigma>1$).
-By Lemma 1.2, the linearization error satisfies $\sum\delta_{n}^{2}/(\gamma_{n}^{(0)})^{2}=O(c^{2})$.
-The spectral identity (Proposition 3.1) therefore holds for $\Xi_{c}$ with $O(c)$ error, giving
-$\log|Z_{c}(\sigma+it)|\leq c\,\zeta_{\alpha}(\sigma)+O(1)<\infty$
-for all $\sigma>1$, $|t|\geq 2$ (using $\log|Z_{0}|=O(1)$ from Remark 3.2).
-Part (e) (Unboundedness on $\sigma=1$).
-The ratio $Z_{c}/Z_{0}$ is a Dirichlet series with generalized Euler product (Remark 6.1), absolutely convergent for $\sigma>1$, hence bounded on $\sigma=2$.
-We show it is unbounded on $\sigma=1$ by a Phragmén–Lindelöf argument (cf.
-[14, Thm.
-8.4]).
+Part (c): the counting function $N_{c}(T)=\#\{n:\gamma_{n}\leq T\}$ satisfies $|N_{c}(T)-N_{\mathrm{main}}(T)|=O(\log T)$.
+Indeed, zeros near height $T$ have index $n\sim T\log T/(2\pi)$ and perturbation $|\delta_{n}|=O(T^{1/2+\varepsilon}/\log T)$ by (2).
+In an interval of this length around $T$, $N_{\mathrm{main}}$ changes by $O(N_{\mathrm{main}}^{\prime}(T)\cdot T^{1/2+\varepsilon}/\log T)=O(T^{1/2+\varepsilon}/\log^{2}T)=o(\log T)$.
+Part (d) (Boundedness for $\sigma>1$). The spectral identity (Proposition 3.1) gives
+$\log|Z_{c}(\sigma+it)|=\operatorname{Re}\log\frac{\Xi_{c}}{\Xi_{0}}(s)+\log|Z_{0}(s)|\leq c\,P(\sigma)+O(1)$
+for all $\sigma>1$, $|t|\geq 2$ (using $|\sum p_{k}^{-\sigma}\cos(\cdots)|\leq P(\sigma)$ and Lemma 3.3).
+Part (e) (Unboundedness on $\sigma=1$). We show $\sup_{t}|\Xi_{c}/\Xi_{0}(1+it)|=+\infty$ by a Phragmén–Lindelöf argument (cf. [17, Thm. 8.4]).
 Step 1.
-$Z_{c}/Z_{0}$ is unbounded in $\{\sigma>1,\;|t|\geq 2\}$.
+$\Xi_{c}/\Xi_{0}$ is unbounded in $\{\sigma>1,\;|t|\geq 2\}$.
 For any $N$, by Dirichlet’s approximation theorem there exists $t\leq(2\pi/\varepsilon)^{N}$ with $|\omega_{k}t+\phi_{k}-\pi|<\varepsilon\pmod{2\pi}$ for all $k\leq N$.
 At $\sigma=1+1/\log N$:
-$\operatorname{Re}\log\frac{Z_{c}}{Z_{0}}(\sigma\!+\!it)\geq c\sum_{k=1}^{N}p_{k}^{-\sigma}-O(c)\geq\tfrac{c}{3}\log\log N-O(1),$
-which grows without bound as $N\to\infty$ (since $\sum_{p\leq N}p^{-1-1/\log N}\geq e^{-2}\sum_{p\leq N}p^{-1}\sim e^{-2}\log\log N$ by Mertens).
+$\operatorname{Re}\log\frac{\Xi_{c}}{\Xi_{0}}(\sigma\!+\!it)\geq c\sum_{k=1}^{N}p_{k}^{-\sigma}-O(c)\geq\tfrac{c}{e}\log\log N-O(1),$
+which grows without bound as $N\to\infty$ (since $p_{k}^{-1/\log N}=e^{-\log p_{k}/\log N}\geq e^{-1}$ for $p_{k}\leq N$, so $\sum_{p\leq N}p^{-1-1/\log N}\geq e^{-1}\sum_{p\leq N}p^{-1}\sim e^{-1}\log\log N$).
 Step 2.
-Suppose $|Z_{c}/Z_{0}(1+it)|\leq M$ for all $t$.
-By Phragmén–Lindelöf in the strip $1\leq\sigma\leq 2$ (the function has at most polynomial growth on $\sigma=1$ and is bounded on $\sigma=2$), $Z_{c}/Z_{0}$ would be bounded throughout the strip.
+Suppose $|\Xi_{c}/\Xi_{0}(1+it)|\leq M$ for all $t$.
+Since $\Xi_{c}$ and $\Xi_{0}$ are entire of order $1$, $|\Xi_{c}(\sigma+it)/\Xi_{0}(\sigma+it)|\leq\exp(C|t|^{1+\varepsilon})$ for any $\varepsilon>0$ (the ratio is analytic in $\sigma\geq 1$ since all zeros of $\Xi_{0}$ lie on $\sigma=1/2$).
+The Phragmén–Lindelöf principle for the strip $1\leq\sigma\leq 2$—bounded on $\sigma=1$ (by assumption) and on $\sigma=2$ (by part (d)), with at most exponential growth—gives boundedness throughout.
 This contradicts Step 1.
-Therefore $Z_{c}/Z_{0}$ is unbounded on $\sigma=1$.
-Since $|Z_{0}(1+it)|=O(1)$ (Remark 3.2) and $|Z_{0}(1+it)|$ is bounded away from zero for $|t|\geq 2$, it follows that $\sup_{t}|Z_{c}(1+it)|=+\infty$.
-The transition is thus: for $\sigma>1$, $|Z_{c}|\leq\exp(c\,\zeta_{\alpha}(\sigma)+O(1))$ (unconditional by Lemma 1.2); on $\sigma=1$, $\sup_{t}|Z_{c}(1+it)|=+\infty$ (unconditional by the Dirichlet–PLF argument above).
-(Note: $Z_{c}$ has a pole at $s=1$ from the factor $1/[\tfrac{1}{2}s(s-1)]$; the large values are for $|t|\to\infty$.) ∎
-6. The Ratio $Z_{c}/Z_{0}$ and Dirichlet Series
-Remark 6.1 (Dirichlet series structure).
-Since $\alpha_{k}=p_{k}$, the spectral identity becomes
-$\log\frac{Z_{c}}{Z_{0}}(s)=-c\sum_{k=1}^{\infty}e^{-i\phi_{k}}\,p_{k}^{-s}+O(c),$
-which is an ordinary Dirichlet series supported on primes. Exponentiating:
-(9) $\frac{Z_{c}}{Z_{0}}(s)=\prod_{p}\exp\!\bigl(-c\,e^{-i\phi_{k}}\,p^{-s}\bigr)=\sum_{n=1}^{\infty}a_{n}\,n^{-s},$
-with $|a_{p}|=c$,   $|a_{p^{m}}|=c^{m}/m!$,   and $a_{mn}=a_{m}a_{n}$ for $\gcd(m,n)=1$.
-This converges absolutely for $\sigma>1$.
-Since both $\Xi_{c}$ and $\Xi_{0}$ satisfy $\Xi(s)=\Xi(1\!-\!s)$ with the same $\Gamma/\pi$ prefactor, the ratio inherits a functional equation:
-$\frac{Z_{c}}{Z_{0}}(s)=\frac{Z_{c}}{Z_{0}}(1-s).$
-The ratio $Z_{c}/Z_{0}$ is meromorphic: it has zeros at $\{1/2+i\gamma_{n}^{(c)}\}$ (from $Z_{c}$) and poles at $\{1/2+i\gamma_{n}^{(0)}\}$ (from $Z_{0}$), both on $\sigma=1/2$.
-It is a Dirichlet series (with generalized Euler product) that is bounded for $\sigma>1$ and unbounded on $\sigma=1$, with all its zeros provably on the critical line.
-The construction separates the Euler product (which lives on the ratio) from the entireness (which lives on $\Xi_{c}$).
-Note that this variant uses $\alpha_{k}=p_{k}$ (rational frequencies) and produces a single function $\Xi_{c}$ with the same properties as in Theorem 1.1; the ratio $Z_{c}/Z_{0}$ is then a Dirichlet series, but $Z_{0}$ is not.
+Therefore $\Xi_{c}/\Xi_{0}$ is unbounded on $\sigma=1$, and since $Z_{c}/Z_{0}=\Xi_{c}/\Xi_{0}$ and $|Z_{0}(1+it)|$ is bounded away from zero for $|t|\geq 2$ (Lemma 3.3), $\sup_{t}|Z_{c}(1+it)|=+\infty$.
+Part (f) (Approximate Euler product). This is Proposition 5.1 below. ∎
+5. The Ratio $\Xi_{c}/\Xi_{0}$ and Dirichlet Series
+Proposition 5.1.
+For $\sigma>1$, the ratio $\Xi_{c}/\Xi_{0}$ admits the approximate Euler product representation
+(12) $\frac{\Xi_{c}}{\Xi_{0}}(s)=\kappa_{0}\cdot\prod_{k=1}^{\infty}\exp\!\bigl(-c\,e^{-ik\theta}\,p_{k}^{-s}\bigr)\cdot M(s),$
+where $\kappa_{0}=e^{cC_{1}}$ is an explicit nonzero constant (with $C_{1}=C_{1}(\theta)$ the non-oscillatory constant from (10)), and $M(s)$ is bounded analytic with $M(s)$ bounded for $\sigma>1/2$, $|t|\geq 2$.
+The Euler product $\prod\exp(-ce^{-ik\theta}p_{k}^{-s})=\sum a_{n}\,n^{-s}$ has $|a_{p_{k}}|=c$, $|a_{p_{k}^{m}}|=c^{m}/m!$, and $a_{mn}=a_{m}a_{n}$ for $\gcd(m,n)=1$.
+Proof.
+The spectral identity (8) gives $\log(\Xi_{c}/\Xi_{0})(s)=-cD(s)+R(s)$ with $R$ bounded.
+Write R(s)=R(2)+(R(s)R(2))R(s)=R(2)+(R(s)-R(2)); the constant $\kappa_{0}=e^{R(2)}$ is a nonzero number and $M(s)=e^{R(s)-R(2)}$ is bounded analytic.
+Then $\Xi_{c}/\Xi_{0}=\kappa_{0}\cdot\exp(-cD(s))\cdot M(s)$.
+Expanding the exponential of the prime-supported Dirichlet series gives the Euler product and multiplicative coefficients.
+∎
+Since both $\Xi_{c}$ and $\Xi_{0}$ satisfy $\Xi(s)=\Xi(1\!-\!s)$, the ratio inherits the functional equation $\Xi_{c}(s)/\Xi_{0}(s)=\Xi_{c}(1-s)/\Xi_{0}(1-s)$.
+The ratio is meromorphic: it has zeros at $\{1/2+i\gamma_{n}\}$ (from $\Xi_{c}$) and poles at $\{1/2+i\gamma_{n}^{(0)}\}$ (from $\Xi_{0}$), all on $\sigma=1/2$.
 6. Almost-Periodicity
-The transition at $\sigma=1$ has a natural interpretation in the hierarchy of almost-periodic function spaces.
+The Dirichlet series $D(s)$ itself has a clean almost-periodicity hierarchy.
 Proposition 6.1.
-Let $\sigma>\tfrac{1}{2}$.
   1. (i)
-    For $\sigma>1$: the function $t\mapsto\log|Z_{c}/Z_{0}|(\sigma\!+\!it)$ is Bohr almost-periodic.
+    For $\sigma>1$: $t\mapsto D(\sigma+it)$ is Bohr almost-periodic (uniformly AP), since $D$ is an absolutely convergent Dirichlet series.
   2. (ii)
-    For $\sigma>\tfrac{1}{2}$: the function $t\mapsto\log|Z_{c}/Z_{0}|(\sigma\!+\!it)$ is Besicovitch ($B^{2}$) almost-periodic with
-    $\|{\log|Z_{c}/Z_{0}|}\|_{B^{2}}^{2}=\tfrac{1}{2}\,c^{2}\sum_{k=1}^{\infty}\alpha_{k}^{-2\sigma}<\infty.$
+    For $\sigma>\tfrac{1}{2}$: $t\mapsto D(\sigma+it)$ is Besicovitch ($B^{2}$) almost-periodic with $\|D(\sigma+i\,\cdot\,)\|_{B^{2}}^{2}=\sum_{k=1}^{\infty}p_{k}^{-2\sigma}<\infty$.
   3. (iii)
-    The construction $E(T)$ is not $B^{2}$-AP: $\sum|a_{k}|^{2}=\frac{1}{\pi^{2}}\sum\alpha_{k}^{-1}=\infty$ (Mertens).
-The truncation $E_{K}\in B^{2}\text{-AP}$ with $\|E_{K}\|_{B^{2}}^{2}=\frac{1}{2\pi^{2}}\sum_{k=1}^{K}\alpha_{k}^{-1}$.
+    The restriction $E(T)=\pi^{-1}\operatorname{Im}\,D(\tfrac{1}{2}+iT)$ is not $B^{2}$-AP: $\sum p_{k}^{-1}=\infty$ (Mertens).
 Proof.
-(i) By Proposition 3.1, the function equals an absolutely convergent trigonometric series $c\sum\alpha_{k}^{-\sigma}\cos(\omega_{k}t+\phi_{k})+O(c)$ for $\sigma>1$.
-Since $\sum\alpha_{k}^{-\sigma}$ converges for $\sigma>1$ [11, Thm.
-2.7], the series defines a Bohr-AP function.
-(ii) The Fourier–Bohr coefficients at exponents $\pm\omega_{k}$ have squared amplitudes $(c\,\alpha_{k}^{-\sigma}/2)^{2}$.
-Since $\sum\alpha_{k}^{-2\sigma}$ converges for $2\sigma>1$, the Riesz–Fischer theorem for $B^{2}$-AP functions gives the result.
-(iii) The amplitudes of $E(T)$ are $a_{k}=1/(\pi\alpha_{k}^{1/2})$.
-Since $\sum|a_{k}|^{2}=(1/\pi^{2})\sum\alpha_{k}^{-1}\to\infty$ (by Mertens’ theorem [11, Thm.
-2.7]), $E$ does not lie in $B^{2}\text{-AP}$ (its mean square diverges).
-For the truncation $E_{K}$, the finite sum confirms $E_{K}\in B^{2}\text{-AP}$.
+(i) Absolute convergence of $\sum p_{k}^{-\sigma}$ for $\sigma>1$ gives uniform convergence of the trigonometric series, hence Bohr-AP.
+(ii) The Fourier–Bohr coefficients of $D(\sigma+it)$ at exponents $\omega_{k}=\log p_{k}$ have squared moduli $p_{k}^{-2\sigma}$.
+Since $\sum p_{k}^{-2\sigma}$ converges for $\sigma>1/2$, the Riesz–Fischer theorem gives the result.
+(iii) Since $\sum p_{k}^{-1}=\infty$ (Mertens), the Fourier energy diverges at $\sigma=1/2$.
 ∎
-The AP hierarchy for $Z_{c}$ mirrors that of $\zeta$: Bohr-AP for $\sigma>1$ (Bohr [1]), $B^{2}$-AP for $\sigma>1/2$ (Bohr–Jessen [3]), with the passage from convergent to divergent Fourier norm at $\sigma=1$ being the “Bohr-AP to $B^{2}$-AP” transition.
-The transitions occur at the same critical $\sigma$ values because $\alpha_{k}\sim p_{k}$; the difference is the source: arithmetic (Euler product) for $\zeta$, purely analytic (spectral identity) for $Z_{c}$.
+The spectral identity transfers these properties to $\log(\Xi_{c}/\Xi_{0})$ up to the bounded error $R(s)$.
+The Fourier–Bohr coefficients of $\log(\Xi_{c}/\Xi_{0})$ at frequencies $\omega_{k}$ are $c\,e^{-ik\theta}\,p_{k}^{-\sigma}$ to leading order; since $R$ is bounded but need not itself be almost-periodic, the AP properties of $\log(\Xi_{c}/\Xi_{0})$ hold to leading order.
+The AP hierarchy mirrors that of $\zeta$: Bohr-AP for $\sigma>1$ (Bohr [1]), $B^{2}$-AP for $\sigma>1/2$ (Bohr–Jessen [3]), with the convergent-to-divergent Fourier norm transition at $\sigma=1$.
 7. Observations
 We illustrate the construction numerically with $c=0.05$ and a finite truncation $K=80$.
-(The computations use the shifted variant $\alpha_{k}=p_{k}+\theta$ with $\theta=e^{\sqrt{2}}$, which ensures global monotonicity of $F_{K}$ (Remark 2.3); the qualitative behavior is identical to the $\alpha_{k}=p_{k}$ case.)
+The computations use the shifted variant $\alpha_{k}=p_{k}+\theta$ in place of $p_{k}$, which ensures global monotonicity of $F_{K}(T)=N_{\mathrm{main}}(T)+cE_{K}(T)$ (see Remark 7.1).
+Since $\alpha_{k}\sim p_{k}$ as $k\to\infty$, the convergence/divergence properties at $\sigma=1$ are unchanged, and the spectral identity has the same structure with $\alpha_{k}$ replacing $p_{k}$.
+Remark 7.1.
+For the finite truncation $E_{K}$, global monotonicity of $F_{K}$ holds when $c<N_{\mathrm{main}}^{\prime}(T_{0})/(\sum_{k=1}^{K}\omega_{k}\alpha_{k}^{-1/2})$, which is satisfied for the parameters used.
+For the full infinite series, occasional crossings ($\gamma_{n+1}<\gamma_{n}$) may occur but do not affect the Hadamard product or the spectral identity (which depend on the linearization at $\gamma_{n}^{(0)}$, not on the ordering of $\gamma_{n}$).
 7.1. First zeros of $Z_{c}$
@@
 Max shift: $|\delta_{n}|\leq 0.077$. Min gap: $1.67>0$. Gaps decrease as $\sim 2\pi/\log(\gamma_{n}/(2\pi))$, matching the Weyl law.
-8.2. Comparison with $\zeta$ zero statistics
-The microscopic statistics of $Z_{c}$’s zeros differ fundamentally from those of $\zeta$. With 1000 zeros computed ($t\in[17.9,1420.1]$):
-Statistic $\zeta$ (conditional on RH) $Z_{c}$ Clock
-RMS of error term $\sqrt{\frac{\log\log T}{2\pi^{2}}}\to\infty$ $0.013$ (constant) 0
-Spacing std (normalized) $\approx 0.42$ (GUE) $0.020$ 0
-Pair correlation $1-\mathrm{sinc}^{2}\alpha$ peaks at integers $\delta$’s
-Number variance $\Sigma^{2}(L)$ $\sim\tfrac{2}{\pi^{2}}\log L$ $\lesssim 0.25$ 0
-Distribution of $S$ / $E$ Gaussian bounded, non-Gaussian 0
-The zeros of $Z_{c}$ are super-rigid (nearly crystalline), with normalized spacing std two orders of magnitude below GUE.
-The construction captures the macroscopic transition (bounded/unbounded, AP hierarchy) but not the microscopic GUE statistics, which arise from the arithmetic content (Euler product, prime correlations) that $Z_{c}$ deliberately lacks.
+7.2. Zero statistics: theory and numerics
+The microscopic statistics of $\Xi_{c}$’s zeros can be analyzed both theoretically and numerically. The key quantity is the ratio of perturbation to spacing:
+(13) $\frac{|\delta_{n}|}{\text{spacing}}\approx c\,|E(\gamma_{n}^{(0)})|.$
+Mean square of $E$ at the zeros. By Weyl equidistribution of the phases $\{\omega_{k}\gamma_{n}^{(0)}+\phi_{k}\}$ modulo $2\pi$ (for each finite truncation $K$, then taking $K\to\infty$):
+(14) $\frac{1}{N}\sum_{n=1}^{N}E(\gamma_{n}^{(0)})^{2}\;\longrightarrow\;\frac{1}{2\pi^{2}}\sum_{k=1}^{\infty}p_{k}^{-1}\;=\;\frac{\log\log N+M}{2\pi^{2}}+o(1)$
+(where $M$ is the Meissel–Mertens constant), so the RMS of $E(\gamma_{n}^{(0)})$ grows like $\sqrt{\log\log N/(2\pi^{2})}$. Consequently:
+  1. (1)
+    RMS of the counting error: $\bigl(\frac{1}{N}\sum(N_{c}(\gamma_{n})-n)^{2}\bigr)^{1/2}\sim c\sqrt{\log\log N/(2\pi^{2})}$, which matches the growth rate of $S(T)$ for $\zeta$ (cf.
+[17, Thm.
+14.24]) but scaled by $c$.
+  2. (2)
+    Perturbation-to-spacing ratio: By (13), the RMS ratio is $c\sqrt{\log\log N/(2\pi^{2})}$, which exceeds $1$ only for $N>\exp\exp(2\pi^{2}/c^{2})$.
+For $c=0.05$ this threshold is $N>10^{10^{347}}$—far beyond any computation.
+Thus the zeros appear nearly crystalline (“super-rigid”) over any accessible range, but asymptotically the perturbations grow to many spacings, and the clock-like character is lost.
+  3. (3)
+    Normalized spacing variance: The deviation of the $n$-th normalized spacing from $1$ is $(\delta_{n+1}-\delta_{n})\cdot N_{\mathrm{main}}^{\prime}=c(E(\gamma_{n+1}^{(0)})-E(\gamma_{n}^{(0)}))$, which has variance $O(c^{2})$ for each $n$ (from the smoothness of $E$ on the scale of zero spacing).
+The normalized spacing std is therefore $O(c)$, much smaller than the GUE value $\approx 0.42$.
+Numerical observations ($c=0.05$, $K=80$, 1000 zeros with $t\in[17.9,1420.1]$):
+Statistic $\zeta$ (conditional on RH) $Z_{c}$ (observed)
+RMS of $N-N_{\mathrm{main}}$ $\sqrt{\frac{\log\log T}{2\pi^{2}}}\to\infty$ $0.013$
+Spacing std (normalized) $\approx 0.42$ (GUE [12]) $0.020$
+Number variance $\Sigma^{2}(L)$ $\sim\tfrac{2}{\pi^{2}}\log L$ $\lesssim 0.25$
+The observed values reflect the regime $c|E|\ll 1$ and are consistent with the theoretical $O(c)$ spacing deviation.
+The asymptotic growth $c\sqrt{\log\log N}$ is invisible at these heights.
+The construction captures the macroscopic transition (bounded/unbounded, AP hierarchy) but not the microscopic GUE statistics of $\zeta$, which arise from the Euler product and prime correlations that $Z_{c}$ deliberately lacks.
 7.3. Value distribution
-Figure 1 shows the complex-plane images of $\log(Z_{c}/Z_{0})(\sigma+it)$, $\log\zeta(\sigma+it)$, and $\log L(\sigma+it,\chi_{4})$ as $t$ ranges over $[20,300]$, for $\sigma=1.25,1.0,0.75$.
-The analytic formula gives
-(10) $\log\frac{Z_{c}}{Z_{0}}(s)=-c\sum_{k=1}^{\infty}\alpha_{k}^{-s}\,e^{i\phi_{k}}+O(c),$
-which has equal real and imaginary RMS: $\langle(\operatorname{Re}\log\frac{Z_{c}}{Z_{0}})^{2}\rangle=\langle(\operatorname{Im}\log\frac{Z_{c}}{Z_{0}})^{2}\rangle=\tfrac{c^{2}}{2}\sum\alpha_{k}^{-2\sigma}$.
+Figure 1 shows the complex-plane images of $\log(\Xi_{c}/\Xi_{0})(\sigma+it)$, $\log\zeta(\sigma+it)$, and $\log L(\sigma+it,\chi_{4})$ as $t$ ranges over $[20,300]$, for $\sigma=1.25,1.0,0.75$.
+By the spectral identity and the $B^{2}$-AP structure of $D$, the mean square of $\log(\Xi_{c}/\Xi_{0})$ satisfies (to leading order in $c$): $\langle(\operatorname{Re}\log\frac{\Xi_{c}}{\Xi_{0}})^{2}\rangle\approx\langle(\operatorname{Im}\log\frac{\Xi_{c}}{\Xi_{0}})^{2}\rangle\approx\tfrac{c^{2}}{2}\sum p_{k}^{-2\sigma}$ (to leading order, from the $B^{2}$-norm of $D$).
 The value distribution is isotropic in the complex plane, as for $\log\zeta$ via Bohr–Jessen [3].
-The quantitative difference is dramatic: at $\sigma=0.75$, the RMS of $\log(Z_{c}/Z_{0})$ is approximately $0.046$, while $\log\zeta$ has RMS $\approx 0.88$.
+The quantitative difference is dramatic: at $\sigma=0.75$, the RMS of $\log(\Xi_{c}/\Xi_{0})$ is approximately $0.046$, while $\log\zeta$ has RMS $\approx 0.88$.
 The construction’s value distribution is confined to a bounded disk, whereas $|\log\zeta(\sigma+it)|$ is unbounded.
-This means $Z_{c}/Z_{0}$ cannot satisfy Voronin universality [15], and its moments do not exhibit the $(\log T)^{k^{2}}$ growth of $\zeta$.
+This means $\Xi_{c}/\Xi_{0}$ cannot satisfy Voronin universality [18], and its moments do not exhibit the $(\log T)^{k^{2}}$ growth of $\zeta$ (cf.
+[11]).
 The value distribution is isotropic but concentrated—another manifestation of the missing arithmetic.
@@
 Complex-plane image of $\log f(\sigma+it)$ for $t\in[20,300]$.
-Left: $\log(Z_{c}/Z_{0})$; center: $\log\zeta$; right: $\log L(s,\chi_{4})$.
+Left: $\log(\Xi_{c}/\Xi_{0})$; center: $\log\zeta$; right: $\log L(s,\chi_{4})$.
 Rows: $\sigma=1.25,1.0,0.75$.
@@
 Figure 2.
-Shape comparison: $\log(Z_{c}/Z_{0})$ rescaled to match the RMS of $\log\zeta$ (left) vs.
+Shape comparison: $\log(\Xi_{c}/\Xi_{0})$ rescaled to match the RMS of $\log\zeta$ (left) vs.
 $\log\zeta$ (right).
-The rescaled $Z_{c}$ fills a banded region (reflecting the coherent trigonometric structure) rather than the isotropic spirals of $\log\zeta$ (reflecting multiplicative independence of primes).
+The rescaled curve fills a banded region (reflecting the coherent trigonometric structure) rather than the isotropic spirals of $\log\zeta$ (reflecting multiplicative independence of primes).
 7.4. Numerical evidence for the linearization condition
-Lemma 1.2 establishes that $\sum_{n=1}^{\infty}E(\gamma_{n}^{(0)})^{2}/n^{2}<\infty$. We verify this numerically with $K=80$ terms and $N=5000$ zero positions.
-Growth of $|E(\gamma_{n}^{(0)})|$.
-The values $|E(\gamma_{n}^{(0)})|$ remain bounded: over the first 5000 zeros, $\max_{n}|E(\gamma_{n}^{(0)})|\approx 0.813$.
-A power-law regression gives $|E(\gamma_{n}^{(0)})|\sim n^{-0.025}$ (essentially flat), and a logarithmic fit yields $|E(\gamma_{n}^{(0)})|\sim(\log n)^{-0.16}$ (very slightly decreasing).
-The Abel-summation bound $|E(T)|=O(T)$ grossly overestimates the actual behavior at zero positions.
+Lemma 3.2 establishes that $\sum_{n=1}^{\infty}E(\gamma_{n}^{(0)})^{2}/n^{2}<\infty$. We verify key aspects numerically with $K=80$ terms and $N=5000$ zero positions.
 Mean square and Weyl equidistribution. The empirical mean square agrees with the Weyl prediction:
-NN $\frac{1}{N}\sum_{n=1}^{N}E(\gamma_{n}^{(0)})^{2}$ $\frac{1}{2\pi^{2}}\sum_{k=1}^{K}\alpha_{k}^{-1}$ ratio
+NN $\frac{1}{N}\sum_{n=1}^{N}E(\gamma_{n}^{(0)})^{2}$ $\frac{1}{2\pi^{2}}\sum_{k=1}^{K}p_{k}^{-1}$ ratio
 100 0.0688 0.0654 1.053
@@
 5000 0.0648 0.0654 0.991
-The ratio stabilizes near $1$, consistent with joint equidistribution of the $K$ phases $\{\omega_{k}\gamma_{n}^{(0)}+\phi_{k}\}$ modulo $2\pi$ (cross terms vanish, diagonal terms give the Mertens sum).
-Convergence of $\sum E^{2}/n^{2}$. The partial sums of $\sum_{n=1}^{N}E(\gamma_{n}^{(0)})^{2}/n^{2}$ stabilize rapidly:
-NN $\sum_{n=1}^{N}E^{2}/n^{2}$ $\sum_{n=1}^{N}\delta_{n}^{2}/\gamma_{n}^{2}$ last term $E^{2}/n^{2}$
-100 0.0891 $3.79\times 10^{-5}$ $2.80\times 10^{-5}$
-1000 0.0897 $88\times 10^{-5} 3.3$$30\times 10^{-7}$
-5000 0.0898 $3.89\times 10^{-5}$ $3.00\times 10^{-11}$
-The growth from $N=1000$ to $N=5000$ is only $0.06\%$. The linearization error $\sum\delta_{n}^{2}/\gamma_{n}^{2}\approx 3.9\times 10^{-5}$ is far below $c^{2}=2.5\times 10^{-3}$.
-Spectral identity remainder. The remainder $R(t)$ in the spectral identity $\log|\Xi_{c}/\Xi_{0}|(\sigma\!+\!it)=-c\sum\alpha_{k}^{-\sigma}\cos(\omega_{k}t+\phi_{k})+R(t)$ has:
+The ratio stabilizes near $1$, consistent with joint equidistribution of the $K$ phases $\{\omega_{k}\gamma_{n}^{(0)}+\phi_{k}\}$ modulo $2\pi$.
+Spectral identity remainder. The remainder $R(t)$ in the spectral identity $\operatorname{Re}\log(\Xi_{c}/\Xi_{0})(\sigma\!+\!it)=-c\sum p_{k}^{-\sigma}\cos(\omega_{k}t+\phi_{k})+R(t)$ has:
 $\sigma$ $\operatorname{mean}(R)$ $\operatorname{std}(R)$ $\max|R|$
@@
 1.00 0.016-0.016 $0.034$ $0.138$
-The remainder is $O(c)$ as claimed in Proposition 3.1, and the contribution from linearization ($\sim 10^{-5}$) is negligible compared to the constant and lower-limit terms.
-9. Open Problems
+The remainder is $O(c)$ as claimed, and the linearization contribution ($\sim 10^{-5}$) is negligible compared to the constant and lower-limit terms.
+8. Open Problems and Concluding Remarks
+The function $\Xi_{c}$ provides an explicit entire function of order $1$ that provably satisfies an analogue of RH (all zeros on $\sigma=1/2$) together with the sharp bounded/unbounded transition at $\sigma=1$ characteristic of $\zeta$.
+The ratio $\Xi_{c}/\Xi_{0}$ is, to leading order, a Dirichlet series with generalized Euler product (Proposition 5.1), providing a concrete object whose zeros are provably on the critical line and whose analytic properties can be studied unconditionally.
+The construction demonstrates that the $\sigma=1$ transition is driven entirely by the marginal convergence/divergence of the Dirichlet series $D(s)=\sum e^{-ik\theta}p_{k}^{-s}$ at $\sigma=1$, in direct analogy with $\log\zeta(s)=P(s)+O(1)$ [17, §3.2].
+In the context of the Selberg class, this shows that an RH-like zero distribution is compatible with the $\sigma=1$ transition—the two properties reinforce rather than obstruct each other.
 Several directions for further investigation suggest themselves.
@@
 The zeros of $Z_{c}$ are super-rigid (nearly crystalline spacing, §7.2), far from the GUE statistics of $\zeta$.
-What additional structure on the zero perturbation $E$ would force GUE pair correlation onto the zeros of $\Xi_{c}$?
+What additional structure on the perturbation would force GUE pair correlation onto the zeros of $\Xi_{c}$?
 Pointwise bounds on $E$.
-Numerically, $|E(\gamma_{n}^{(0)})|\leq 1.05$ for all computed nn8.4), vastly smaller than the Abel bound $O(\gamma_{n})$.
-Proving $|E(\gamma_{n}^{(0)})|=O(1)$ would give a much tighter linearization estimate than Lemma 1.2.
+The mean square (14) shows that $|E(\gamma_{n}^{(0)})|\sim\sqrt{\log\log n}$ in RMS, but the worst-case bound (2) gives only $O(\gamma_{n}^{1/2+\varepsilon})$.
+Numerically, $|E(\gamma_{n}^{(0)})|\leq 1.05$ for all computed nn7.4).
+Proving that $|E(\gamma_{n}^{(0)})|=O(\sqrt{\log\log n})$ pointwise (matching the RMS rate) would sharpen the linearization estimate and clarify when the perturbations first exceed the zero spacing.
 Growth rate on $\sigma=1$.
 The Phragmén–Lindelöf argument gives $\sup_{t}|Z_{c}(1+it)|=+\infty$ but not the rate.
-For $\zeta$, $|\zeta(1+it)|=\Omega(\log\log t)$ [14, Thm.
+For $\zeta$, $|\zeta(1+it)|=\Omega(\log\log t)$ [17, Thm.
 8.5].
 Does $|Z_{c}(1+it)|$ achieve the same rate?
-10. Concluding Remarks
-The function $\Xi_{c}$ provides the first explicit entire function of order $1$ that provably satisfies an analogue of the Riemann Hypothesis (all zeros on $\sigma=1/2$) together with the sharp bounded/unbounded transition at $\sigma=1$ characteristic of $\zeta$.
-The construction demonstrates that these properties are simultaneously achievable and provable, given sufficient control over the zero distribution.
-The ratio $Z_{c}/Z_{0}$ is a Dirichlet series with generalized Euler product, providing a concrete object whose zeros are provably on the critical line and whose analytic properties (boundedness, value distribution, almost-periodicity) can be studied unconditionally.
-In the context of the Selberg class [12, 9], the construction shows that an RH-like zero distribution is compatible with the $\sigma=1$ transition—the two properties reinforce rather than obstruct each other.
-The microscopic statistics of $Z_{c}$’s zeros differ from those of $\zeta$8.2): the zeros are nearly crystalline (super-rigid spacing) rather than GUE-distributed, and $Z_{c}$ does not satisfy Voronin universality.
-These differences reflect the specific arithmetic content—prime correlations, multiplicative independence of $p^{-it}$—that $\zeta$ possesses and $Z_{c}$ deliberately lacks.
-Understanding which additional structure would force GUE statistics onto the zeros of an RH-satisfying function remains an interesting open problem.
+An actual Dirichlet series.
+The ratio $\Xi_{c}/\Xi_{0}$ is approximately a Dirichlet series with Euler product (Proposition 5.1), but the bounded correction factor $M(s)$ prevents it from being one exactly.
+The base function $\Xi_{0}$ is not a Dirichlet series either (its Fourier exponents $\log\gamma_{n}^{(0)}$ are not of the form $\log n$).
+Can one unconditionally construct a Dirichlet series $\sum a_{n}\,n^{-s}$ (convergent for $\sigma>1/2$) with all zeros provably on $\sigma=1/2$ and the $\sigma=1$ transition?
+The Davenport–Heilbronn example [4] shows that a functional equation alone does not suffice; some form of multiplicative structure appears necessary.
+Acknowledgments
+This paper was developed through a collaboration between the author and Claude (Anthropic).
+The author posed the original question (whether one can construct an explicit entire function with the Riemann Hypothesis and the $\sigma=1$ transition), proposed the unit-disc analogy that motivated the Hadamard product approach, identified that a family result is insufficient (insisting on a single function), and formulated the equidistribution criterion for unboundedness.
+Claude proposed the Hadamard product construction, developed the spectral identity and its proof, found the dyadic large-sieve argument that resolved the linearization convergence, suggested the $\alpha_{k}=p_{k}$ switch that yields the Dirichlet series structure, and wrote the paper.